### v1

# references.bib

Abstract
We compute the Hilbert series of the space of $n=3$ variable quasi-invariant polynomials in characteristic $2$ and $3,$ capturing the dimension of the homogeneous components of the space, and explicitly describe the generators in the characteristic $2$ case. In doing so we extend the work of the first author in 2023 on quasi-invariant polynomials in characteristic $p>n$ and prove that a sufficient condition found by Ren–Xu in 2020 on when the Hilbert series differs between characteristic 00 and $p$ is also necessary for $n=3,p=2,3$. This is the first description of quasi-invariant polynomials in the case where the space forms a modular representation over the symmetric group, bringing us closer to describing the quasi-invariant polynomials in all characteristics and numbers of variables.
1 Introduction
Let $\mathbbm{k}$ be a field, and consider the action of the symmetric group $S_{n}$ on the space $\mathbbm{k}[x_{1},\dots,x_{n}]$ of $\mathbbm{k}$-valued polynomials by permuting the variables. A polynomial in $\mathbbm{k}[x_{1},\dots,x_{n}]$ is symmetric if it is invariant under this action. Equivalently, since $S_{n}$ is generated by transpositions, a polynomial $K$ is symmetric if $s_{ij}K=K$ or $(1-s_{ij})K=0$ for all $s_{ij}\in S_{n}$. One may consider generalizations of symmetric polynomials in which this condition is relaxed, so that we only require $(1-s_{ij})K$ be divisible by some large polynomial. This leads to the notion of quasi-invariant polynomials:
Definition 1.1.
Let $\mathbbm{k}$ be a field. For $m\in\mathbb{Z}_{\geq 0},n\in\mathbb{Z}_{>0}$, a polynomial $K\in\mathbbm{k}[x_{1},\dots,x_{n}]$ is $m$-quasi-invariant if for all $s_{ij}\in S_{n}$ we have that $(x_{i}-x_{j})^{2m+1}$ divides $(1-s_{ij})K$. We denote the space of $m$-quasi-invariants by $Q_{m}(n,\mathbbm{k})$.
Note that the symmetric polynomials are exactly the polynomials that are $m$-quasi-invariant for all $m$. For brevity, we also refer to quasi-invariant polynomials as simply quasi-invariants.
Quasi-invariant polynomials were first introduced by Chalykh and Veselov in 1990 [cmp/1104179957] to describe the harmonic, zero eigenvalue eigenfunctions of quantum Calogero-Moser systems. Calogero-Moser systems are a collection of one-dimensional dynamical particle systems that were found to be both solvable [calogero] and integrable [MOSER1975197]. Due to these properties, they have become extensively studied in mathematical physics, with connections to a number of other fields of mathematics, including representation theory.
Quasi-invariant polynomials were also later found to describe the representation theory of the spherical subalgebra of the rational Cherednik algebra [beg]. This subalgebra is Morita equivalent to the entire rational Cherednik algebra [Etingof2000SymplecticRA], so quasi-invariants describe representations of rational Cherednik algebras as well. Such algebras have connections to combinatorics, mathematical physics, algebraic geometry, algebraic topology, and more, leading them to become a central topic in representation theory.
Due to these applications, the quasi-invariant polynomials have been studied extensively in recent years. Of particular interest are properties such as its freeness as a module over the symmetric polynomials and the degrees of its generators. To describe these properties, it is useful to consider the Hilbert series of the quasi-invariants, which encapsulates much of this information.
Definition 1.2.
Let $V=\bigoplus_{d=0}^{\infty}V_{d}$ be a graded vector space. The Hilbert series of $V$ is the formal power series
$\mathcal{H}(V):=\sum_{d=0}^{\infty}\dim(V_{d})t^{d}.$
In 2003, Felder and Veselov found the Hilbert series of the space of quasi-invariants in characteristic zero [felder2001action], proving its freeness in the process. Work on quasi-invariants in characteristic $p$ started in 2020, when Ren and Xu proved a sufficient condition for the Hilbert series of $Q_{m}(n,\mathbf{F}_{p})$ to be different from the Hilbert series of $Q_{m}(n,\mathbf{Q})$ [Ren_2020]. They accomplished this by computing non-symmetric polynomial “counterexamples” in characteristic $p$ where the polynomial has lower degree than any non-symmetric quasi-invariant polynomial in characteristic 0. They also made several conjectures about quasi-invariants in characteristic $p$, including that the condition they found is also sufficient, the quasi-invariants are free, and that the Hilbert polynomial is palindromic for $p>2$. In 2023, the first author proved a general form for the Hilbert series of the quasi-invariants for $n=3,p>3$, proving freeness and the palindromicity of the Hilbert polynomial in the process [wang2022explicit].
In this paper, we consider the cases $n=3,p=2,3$. These cases differ from the $p>3$ case studied in [wang2022explicit] since in $p=2,3$ the representations of $S_{3}$ are modular, i.e. are not completely reducible. Despite these limitations, we describe the Hilbert series explicitly for all $m$, proving the following:
Theorem 1.3.
Let $\mathbbm{k}$ be either $\mathbf{F}_{2}$ or $\mathbf{F}_{3}$. Then the Hilbert series for $Q_{m}(3,\mathbbm{k})$ is given by
$\mathcal{H}(Q_{m}(3,\mathbbm{k}))=\frac{1+2t^{d}+2t^{6m+3-d}+t^{6m+3}}{(1-t)(1-t^{2})(1-t^{3})}$
where $d=3m+1$ if there is no Ren–Xu counterexample and $d$ is the degree of the minimal degree Ren–Xu counterexample otherwise. In particular, the conditions found in [Ren_2020] for the Hilbert series of $Q_{m}(3,\mathbbm{k})$ to be different from the Hilbert series of $Q_{m}(3,\mathbf{Q})$ are necessary.
Note that this result also implies freeness and the palindromicity of the Hilbert polynomial.
In the case $p=2$, we also define $m$-quasi-invariants in the case where $m$ is a half-integer and prove an analogous statement to Theorem 1.3 in this case. Using quasi-invariants at half-integers, we also compute the generators of $Q_{m}(3,\mathbf{F}_{2})$ as a $\mathbf{F}_{2}[x_{1},x_{2},x_{3}]^{S_{3}}$-module explicitly.
In Section 2, we state some of the basic facts about quasi-invariant polynomials and introduce modular representations of $S_{3}$. In Section 3, we compute the generators of $Q_{m}(3,\mathbf{F}_{2})$, proving Theorem 1.3 for $p=2$ in the process. In Section 4, we begin discussing $p=3$, and show that some properties of quasi-invariants in 3 variables from [wang2022explicit] carry over to the $p=3$ case after converting from the standard representation to the $\mathrm{sign-triv}$ representation. In Section 5, we show that minimal degree Ren–Xu counterexamples are the lowest degree non-symmetric generators for $Q_{m}(3,\mathbf{F}_{3})$ and show that there is one other higher degree generator belonging to the $\mathrm{sign-triv}$ representation. Finally, in Section 6 we consider all other indecomposable representations of $S_{3}$ in $Q_{m}(3,\mathbf{F}_{3})$, finishing the proof of Theorem 1.3 for $p=3$.
2 Preliminaries
We start with some useful properties of the quasi-invariants.
Proposition 2.1 ([etingof2002lectures]).
Let $\mathbbm{k}$ be a field.
  1. 1)
    $\mathbbm{k}[x_{1},x_{2},x_{3}]^{S_{3}}\subset Q_{m}(3,\mathbbm{k})$, $Q_{0}(3,\mathbbm{k})=\mathbbm{k}[x_{1},x_{2},x_{3}],$ and $Q_{m}(3,\mathbbm{k})\supset Q_{m^{\prime}}(3,\mathbbm{k})$ where $m^{\prime}>m$.
  2. 2)
    $Q_{m}(3,\mathbbm{k})$ is a ring.
  3. 3)
    $Q_{m}(3,\mathbbm{k})$ is a finitely generated $\mathbbm{k}[x_{1},x_{2},x_{3}]^{S_{3}}$-module.
We consider $Q_{m}(3,\mathbf{F}_{2})$ and $Q_{m}(3,\mathbf{F}_{3})$ as representations of $S_{3}$ where $S_{3}$ permutes the variables $x_{1},x_{2},x_{3}$. Since $Q_{m}(3,\mathbf{F}_{2})$ and $Q_{m}(3,\mathbf{F}_{3})$ are vector spaces over $\mathbf{F}_{2}$ and $\mathbf{F}_{3}$ respectively and the characteristics $2$ and $3$ divide $|S_{3}|$, $Q_{m}(3,\mathbf{F}_{2})$ and $Q_{m}(3,\mathbf{F}_{3})$ are modular representations of $S_{3}$.
Proposition 2.2.
$Q_{m}(3,\mathbf{F}_{2})$ and $Q_{m}(3,\mathbf{F}_{3})$ are modular representations of $S_{3}$.
First, we consider characteristic $2$.
2.1 Preliminary definitions for p=2
We describe the indecomposable and irreducible representations of $S_{3}$ for $p=2$.
Proposition 2.3 ([Alperin_1986]).
There are $3$ irreducible or indecomposable representations of $S_{3}$ in characteristic $2$:
  1. 1)
    $\mathrm{triv}$ is the irreducible representation of $S_{3}$ that is acted on trivially by $S_{3}.$
  2. 2)
    $\mathrm{std}$ is the $2$ dimensional irreducible representation of $S_{3}$ obtained by reducing the standard representation in characteristic 0 mod 2.
  3. 3)
    $\mathrm{triv-triv}$ is the $2$ dimensional indecomposable representation that contains a copy of $\mathrm{triv}$ as a subrepresentation such that the quotient of $\mathrm{triv-triv}$ by this subrepresentation is $\mathrm{triv}$.
Example 2.4.
The polynomial $E_{\mathrm{triv-triv}}:=x_{1}^{2}x_{2}+x_{2}^{2}x_{3}+x_{3}^{2}x_{1}\in\mathbf{F}_{2}[x_{1},x_{2},x_{3}]$ generates a copy of $\mathrm{triv-triv}$. To see this, note that for any $i,j$, we have
$(1-s_{ij})E_{\mathrm{triv-triv}}=x_{1}^{2}x_{2}+x_{1}x_{2}^{2}+x_{1}^{2}x_{3}+x_{1}x_{3}^{2}+x_{2}^{2}x_{3}+x_{2}x_{3}^{2}\in\mathbf{F}_{2}[x_{1},x_{2},x_{3}]^{S_{3}}.$
Since the transpositions generate $S_{3}$, $E_{\mathrm{triv-triv}}$ generates a two-dimensional representation that contains $\mathrm{triv}$ as a subrepresentation. Moreover, since $E_{\mathrm{triv-triv}}$ is not symmetric, this representation is not $\mathrm{triv}\oplus\mathrm{triv}$, so it must be $\mathrm{triv-triv}$.
We then study the behaviors of each indecomposable representation in the quasi-invariants. We define $Q_{m}(3,\mathbf{F}_{2})_{\mathrm{triv}}$ and $Q_{m}(3,\mathbf{F}_{2})_{\mathrm{std}}$ to be the direct sum of all copies of $\mathrm{triv}$ and $\mathrm{std}$ respectively in the quasi-invariants. We also define $Q_{m}(3,\mathbf{F}_{2})_{\mathrm{triv-triv}}$ to be the direct sum of all copies of $\mathrm{triv}$ and $\mathrm{triv-triv}$.
Remark 2.5.
We cannot define $Q_{m}(3,\mathbf{F}_{2})_{\mathrm{triv-triv}}$ to exclude copies of $\mathrm{triv}$ since we can add elements of $Q_{m}(3,\mathbf{F}_{2})_{\mathrm{triv}}$ to copies of $\mathrm{triv-triv}$ and still obtain a copy of $\mathrm{triv-triv}.$ For example, $F:=E_{\mathrm{triv-triv}}+x_{1}^{3}+x_{2}^{3}+x_{3}^{3}$ still satisfies $(1-s_{ij})F=(1-s_{ij})E_{\mathrm{triv-triv}}$ for all $i,j$, so it generates a copy of $\mathrm{triv-triv}$ by the same argument as Example 2.4.
Proposition 2.6.
[wang2022explicit] As an $\mathbf{F}_{2}[x_{1},x_{2},x_{3}]^{S_{3}}$-module, $Q_{m}(3,\mathbf{F}_{2})_{\mathrm{triv}}$ is freely generated by $1$
Note that by the classification of indecomposables in Proposition 2.3, every extension of $\mathrm{std}$ and every extension of a module by $\mathrm{std}$ splits. Thus $Q_{m}(3,\mathbf{F}_{2})_{\mathrm{std}}$ is a direct summand of $Q_{m}(3,\mathbf{F}_{2})$ (whose complement is $Q_{m}(3,\mathbf{F}_{2})_{\mathrm{triv-triv}}$), and we mainly consider $Q_{m}(3,\mathbf{F}_{2})_{\mathrm{std}}$. $Q_{m}(3,\mathbf{F}_{2})_{\mathrm{std}}$ is generated as a $\mathbf{F}_{2}[x_{1},x_{2},x_{3}]^{S_{3}}$-module by homogeneous copies of $\mathrm{std}$, so following [wang2022explicit], we consider generating representations of $Q_{m}(3,\mathbf{F}_{2})_{\mathrm{std}}$ as homogeneous copies of $\mathrm{std}$ in a generators and relations presentation of $Q_{m}(3,\mathbf{F}_{2})_{\mathrm{std}}$ with a minimal generator set.
2.1.1 Quasi-invariants at half-integers
Note that if $\mathbbm{k}$ is a field with char $\text{char\,}\mathbbm{k}\neq 2$ and $m\in\mathbb{Z}_{\geq 0}$, then for any $K\in\mathbbm{k}[x_{1},\dots,x_{n}]$, $(x_{i}-x_{j})^{2m}|(1-s_{ij})K$ implies $(x_{i}-x_{j})^{2m+1}|(1-s_{ij})K$ since $(1-s_{ij})K$ is $s_{ij}$-antiinvariant, hence the exponent $12m+$ 1 in the definition of quasi-invariant polynomials. But this does not hold in characteristic 2, since there is no concept of antiinvariants. Indeed, one can check that for $K=x_{1}^{2}+x_{2}^{2}$, we have $(x_{i}-x_{j})^{2}|(1-s_{ij})K$ for all $i,j$, but $(x_{i}-x_{j})^{3}\nmid|(1-s_{ij})K$ if $i=1,2$, $j\neq 1,2$.
We encapsulate this data by extending the definition of quasi-invariants to half-integers when $p=2$. For example, $K=x_{1}^{2}+x_{2}^{2}$ is $\frac{1}{2}$-quasi-invariant, and this is in fact the minimal degree nonsymmetric $\frac{1}{2}$-quasi-invariant polynomial. Proposition 2.1 still holds when $m,m^{\prime}$ are half-integers, and the definitions of $Q_{m}(3,\mathbf{F}_{2})_{\mathrm{triv}}$, $Q_{m}(3,\mathbf{F}_{2})_{\mathrm{std}}$ also naturally extend to half-integer $m$. So from now on, whenever we refer to quasi-invariants in characteristic $2$ we let $m$ be a half-integer.
2.2 Preliminary definitions for $p=3$
Next, we define the indecomposable and irreducible representations of $S_{3}$.
Proposition 2.7 ([Alperin_1986]).
There are $6$ indecomposable or irreducible representations in $S_{3}$ in characteristic $3$:
  1. 1)
    $\mathrm{triv}$ is the irreducible representation of $S_{3}$ that is acted on trivially by $S_{3}$.
  2. 2)
    $\mathrm{sign}$ is the irreducible representation of $S_{3}$ that is acted on by negation by the transpositions.
  3. 3)
    $\mathrm{sign-triv}$ is the indecomposable representation that contains a copy of $\mathrm{triv}$ as a subrepresentation, such that the quotient of $\mathrm{sign-triv}$ by this subrepresentation is $\mathrm{sign}$.
  4. 4)
    $\mathrm{triv-sign}$ is the indecomposable representation that contains a copy of $\mathrm{sign}$ as a subrepresentation, such that the quotient of $\mathrm{triv-sign}$ by this subrepresentation is $\mathrm{triv}$.
  5. 5)
    $\mathrm{triv-sign-triv}$ is the indecomposable representation that contains a copy of $\mathrm{sign-triv}$ as a subrepresentation, such that the quotient of $\mathrm{triv-sign-triv}$ by this subrepresentation is $\mathrm{triv}$.
  6. 6)
    $\mathrm{sign-triv-sign}$ is the indecomposable representation that contains a copy of $\mathrm{triv-sign}$ as a subrepresentation, such that the quotient of $\mathrm{sign-triv-sign}$ by this subrepresentation is $\mathrm{sign}$.
Example 2.8.
The space $V\subset\mathbf{F}_{3}[x_{1},x_{2},x_{3}]$ consisting of homogeneous linear polynomials is a copy of $\mathrm{triv-sign-triv}$. Indeed, the space $T\subset V$ spanned by $x_{1}+x_{2}+x_{3}$ is a copy of $\mathrm{triv}$. One can check $x_{1}-x_{2}\in V/T$ is acted by negation by all transpositions in $S_{3},$ so $x_{1}-x_{2}$ generates a copy of $\mathrm{sign}$ in $V/T$. Let $W$ be spanned by $x_{1}-x_{2}$ and $x_{1}+x_{2}+x_{3}$. Then $V/W$ is one dimensional, and one can check that it is a copy of $\mathrm{triv}$. Finally, it is easy to show that there are no copies of $\mathrm{triv}$ or $\mathrm{sign}$ in $V$ other than $T$. In particular, since $V$ has a unique irreducible subrepresentation, it is indecomposable, and we conclude that it is a copy of $\mathrm{triv-sign-triv}$.
Similarly to the $p=2$ case, we define $Q_{m}(3,\mathbf{F}_{3})_{\mathrm{sign}}$ and $Q_{m}(3,\mathbf{F}_{3})_{\mathrm{triv}}$ to be the direct sum of all copies of $\mathrm{sign}$ and $\mathrm{triv}$ in $Q_{m}(3,\mathbf{F}_{3})$ respectively.
Proposition 2.9 ([wang2022explicit]).
As $\mathbf{F}_{3}[x_{1},x_{2},x_{3}]^{S_{3}}$-modules,
  1. 1)
    $Q_{m}(3,\mathbf{F}_{3})_{\mathrm{triv}}$ is freely generated by $1$.
  2. 2)
    $Q_{m}(3,\mathbf{F}_{3})_{\mathrm{sign}}$ is freely generated by $\prod_{i<j}(x_{i}-x_{j})^{2m+1}$.
Next we define $Q_{m}(3,\mathbf{F}_{3})_{\mathrm{sign-triv}}$ as the direct sum of all copies of $\mathrm{sign},$ $\mathrm{triv},$ and $\mathrm{sign-triv}.$ For this paper we consider generators of $Q_{m}(3,\mathbf{F}_{3})_{\mathrm{sign-triv}}$ to be homogeneous polynomials other than $1$ and $\prod_{i<j}(x_{i}-x_{j})^{2m+1}$ such that they are in the $(-1)$-eigenspace of $s_{12}$ and are in a generators and relations presentation of $Q_{m}(3,\mathbf{F}_{3})_{\mathrm{sign-triv}}$ as an $F_{3}[x_{1},x_{2},x_{3}]-$module with the least number of generators. Moreover, if $K$ is a generator of $Q_{m}(3,\mathbf{F}_{3})_{\mathrm{sign-triv}}$ then it necessarily generates a copy of $\mathrm{sign-triv}$ since we assumed $K$ neither generates $\mathrm{triv}$ nor $\mathrm{sign}.$
Remark 2.10.
Similar to in the $p=2$ case, we cannot define $Q_{m}(3,\mathbf{F}_{3})_{\mathrm{sign-triv}}$ to exclude copies of $\mathrm{sign}$ since we can add elements of $Q_{m}(3,\mathbf{F}_{3})_{\mathrm{sign}}$ to copies of $\mathrm{sign-triv}$ and still obtain a copy of $\mathrm{sign-triv}.$ For example, the spaces spanned by $(x_{1}^{6}-x_{2}^{6})(x_{1}+x_{2}+x_{3})^{3}$, $(x_{1}^{6}+x_{2}^{6}+x_{3}^{6})(x_{1}+x_{2}+x_{3})^{3}$ and $\prod_{i<j}(x_{i}-x_{j})^{3}+(x_{1}^{6}-x_{2}^{6})(x_{1}+x_{2}+x_{3})^{3}$, $(x_{1}^{6}+x_{2}^{6}+x_{3}^{6})(x_{1}+x_{2}+x_{3})^{3}$ generate two copies of $\mathrm{sign-triv}$ in $Q_{1}(3,\mathbf{F}_{3})$, and their sum contains $\prod_{i<j}(x_{i}-x_{j})^{3}\in Q_{1}(3,\mathbf{F}_{3})_{\mathrm{sign}}$.
Remark 2.11.
One could define subspaces of $Q_{m}(3,\mathbf{F}_{3})$ for $\mathrm{triv-sign-triv},\mathrm{sign-triv-sign},$ $\mathrm{triv-sign}$ similar to $Q_{m}(3,\mathbf{F}_{3})_{\mathrm{sign-triv}}$, however this is not particularly helpful, as unlike in $p=2$, we cannot decompose $Q_{m}(3,\mathbf{F}_{3})$ into a direct sum of subspaces of this form. $Q_{m}(3,\mathbf{F}_{3})_{\mathrm{sign-triv}}$ is still relevant, as it is the critical piece to understanding quasi-invariants in characteristic 3, as we see in Sections 4 and 5.
3 Quasi-invariants in characteristic 2
In this section we write down explicit generators for $Q_{m}(3,\mathbf{F}_{2})$ and prove Theorem 1.3 for $p=2$. Note that we already know the structure of $Q_{m}(3,\mathbf{F}_{2})_{\mathrm{triv}}$ from Proposition 2.6. We start by extending this to $Q_{m}(3,\mathbf{F}_{2})_{\mathrm{triv-triv}}$.
Proposition 3.1.
As a $\mathbf{F}_{2}[x_{1},x_{2},x_{3}]^{S_{3}}$-module, $Q_{m}(3,\mathbf{F}_{2})_{\mathrm{triv}-\mathrm{triv}}$ is freely generated by $1$ and $E_{\mathrm{triv-triv}}\prod(x_{i}-x_{j})^{2m}$.
Proof.
Let $K$ be a nonsymmetric element of $Q_{m}(3,\mathbf{F}_{2})_{\mathrm{triv}-\mathrm{triv}}$ so that $(x_{i}+x_{j})^{2m+1}|(1+s_{ij})K.$ Because
$(1+s_{12})K=(1+s_{13})K=(1+s_{23})K,$
$(1+s_{ij})K=P\prod(x_{i}-x_{j})^{2m+1}$ for some symmetric polynomial $P.$ Letting $G=E_{\mathrm{triv-triv}}\prod(x_{i}-x_{j})^{2m}$ yields $(1+s_{ij})G=\prod(x_{i}-x_{j})^{2m+1}.$ Thus $(1+s_{ij})PG=(1+s_{ij})K$ and $(1+s_{ij})(PG-K)=0,$ so $PG-K$ is symmetric and $K$ is generated by $G$ and $1$ Moreover since $G$ is not symmetric, $P$ and $G$ have no relation implying freeness. ∎
We have an explicit description of $Q_{m}(3,\mathbf{F}_{2})_{\mathrm{triv-triv}}$, so it remains to compute the generators and relations of $Q_{m}(3,\mathbf{F}_{2})_{\mathrm{std}}$. A number of the properties of $Q_{m}(3,\mathbf{F}_{p})$ for $p>3$ found in [wang2022explicit] are true for $Q_{m}(3,\mathbf{F}_{2}).$ We prove these first.
If $V$ is a copy of $\mathrm{std}$, then we denote by $V_{ij}$ the $1$-eigenspace of $s_{ij}$ in $V$.
Lemma 3.2.
[…]
$(x_{1}-x_{2})^{2m+1}K^{\prime}+(x_{2}-x_{3})^{2m+1}sK^{\prime}+(x_{3}-x_{1})^{2m+1}s^{2}K^{\prime}=0.$
Then $(x_{1}-x_{2})^{2m+1}K^{\prime}$ belongs to the $1$-eigenspace of $s_{12}$ in some copy of $\mathrm{std}$ inside $Q_{m}(3,\mathbf{F}_{2})_{\mathrm{std}}$.
Proof.
For the first statement, $K+sK+s^{2}K=0$ holds for any copy of $\mathrm{std}.$ For the next, suppose $\{i,j,l\}=\{1,2,3\}$ for some integer $l.$ Then $(1-s_{il})K=s_{jl}K,$ so $(x_{i}-x_{l})^{2m+1}|s_{jl}K$, implying $(x_{i}-x_{j})^{2m+1}|K.$ The second statement follows from the proof in [wang2022explicit]. ∎
Corollary 3.3.
Let $V$ be a generating representation of $Q_{m}(3,\mathbbm{k})_{\mathrm{std}}$ and let $K\in V_{ij}$. Let us write $K=(x_{i}-x_{j})^{2m+1}K^{\prime}$. Then $K^{\prime}$ is not divisible by any nonconstant symmetric polynomial.
The proof of this statement is identical to the one in [wang2022explicit].
Lemma 3.4.
Let $V,W$ be distinct generating representations of $Q_{m}(3,\mathbbm{k})_{\mathrm{std}}$. Let $K\in V_{12},L\in W_{12}$. Then $KL+s_{13}Ks_{23}L$ is a nonsymmetric element of $Q_{m}(3,\mathbbm{k})_{\mathrm{triv-triv}}$ and we have $\deg V+\deg W\geq 6m+3$.
Proof.
$KL+s_{13}Ks_{23}L$ is an element of $Q_{m}(3,\mathbf{F}_{2})$ since the quasi-invariants form a ring by Proposition 2.1. We have that
$(1+s_{12})(KL+s_{13}Ks_{23}L)=s_{23}Ks_{13}L+s_{13}Ks_{23}L,$
$(1+s_{13})(KL+s_{13}Ks_{23}L)=KL+s_{13}Ks_{23}L+s_{13}Ks_{13}L+Ks_{23}L=Ks_{13}L+s_{13}KL,$
and
$(1+s_{23})(KL+s_{13}Ks_{23}L)=KL+s_{13}Ks_{23}L+s_{23}Ks_{23}L+s_{13}KL=Ks_{23}L+s_{23}KL.$
Using that $K$ and $L$ lie in copies of $\mathrm{std},$ one can check that these are each the same symmetric polynomial. Thus $KL+s_{13}Ks_{23}L$ lies in a quotient of a copy of $\mathrm{triv-triv}$. Note that by the same argument as in [wang2022explicit], we have $Ks_{23}L+s_{23}KL\neq 0$, so $KL+s_{13}Ks_{23}L$ is nonsymmetric and must generate a copy of $\mathrm{triv-triv}$.
By Proposition 2.6, $KL+s_{13}Ks_{23}L$ has degree at least $6m+3,$ so $\deg V+\deg W\geq 6m+3$ as desired. ∎
Lemma 3.5.
Assume that there exist generating representations $V,W$ of $Q_{m}(3,\mathbf{F}_{2})_{\mathrm{std}}$ such that $\deg V+\deg W=6m+3$. Then $Q_{m}(3,\mathbf{F}_{2})_{\mathrm{std}}$ is a free module over $\mathbbm{k}[x_{1},x_{2},x_{3}]^{S_{3}}$ generated by $V$ and $W$.
Proof.
Assume for the sake of contradiction there exists another generator $U$ of $Q_{m}(3,\mathbf{F}_{2})_{\mathrm{std}}.$ Supposing $\deg W\geq\deg V$, by Lemma 3.4, $\deg U\geq\deg W.$ By Lemma 3.4, if $K\in V_{12},$ $L\in W_{12},$ and $T\in U_{12}$ then $KL+s_{13}Ks_{23}L$ and $KT+s_{13}Ks_{23}T$ are both nonsymmetric elements of $Q_{m}(3,\mathbf{F}_{2})_{\mathrm{triv-triv}}.$ Moreover, we have
$(1+s_{12})(KL+s_{13}Ks_{23}L)=s_{23}Ks_{13}L+s_{13}Ks_{23}L=\prod(x_{i}-x_{j})^{2m+1},$
and
$(1+s_{12})(KT+s_{13}Ks_{23}T)=s_{23}Ks_{13}T+s_{13}Ks_{23}T=Q\prod(x_{i}-x_{j})^{2m+1}$
for some symmetric polynomial $Q.$ From there we may proceed identically to [wang2022explicit]. ∎
Now, we are ready to prove Theorem 1.3 for $p=2$.
Theorem 3.6.
Let $a$ be the largest natural number such that $12^{a}<2m+$ 1 . Then $Q_{m}(3,\mathbf{F}_{2})_{\mathrm{std}}$ is freely generated by $(x_{1}-x_{2})^{2^{a+1}}$ and $(x_{1}-x_{2})^{2^{a}}\prod(x_{i}-x_{j})^{2m+1-2^{a}}.$
Remark 3.7.
Note that when $m$ is an integer, the degrees of the generators in this theorem agree with the degrees conjectured in [Ren_2020]. In particular, when $2^{a+1}$ is one of $23m+1,3m+$ 2, we actually have that the Hilbert series of $Q_{m}(3,\mathbf{F}_{2})$ and $Q_{m}(3,\mathbf{Q})$ agree, so $(x_{1}-x_{2})^{2^{a+1}}$, $(x_{1}-x_{2})^{2^{a}}\prod(x_{i}-x_{j})^{2m+1-2^{a}}$ are the reductions modulo 2 of the generators of $Q_{m}(3,\mathbf{Q})$, when written as integer polynomials with coprime coefficients.
Proof of Theorem 3.6.
We prove this by induction on $m.$
The generators of $Q_{0}(3,\mathbf{F}_{2})_{\mathrm{std}}$ are $(x_{1}-x_{2})$ and $(x_{1}-x_{2})^{2}$, completing our base case.
Let $j$ be a half-integer, and suppose that $Q_{j-\frac{1}{2}}(3,\mathbf{F}_{2})_{\mathrm{std}}$ is freely generated by $(x_{1}-x_{2})^{2^{a+1}}$ and $(x_{1}-x_{2})^{2^{a}}\prod(x_{i}-x_{j})^{2j-2^{a}}$ where $2^{a}$ is the greatest such power of $2$ less than $2j.$ If $2j\neq 2^{a+1},$ then $2^{a}$ is the largest power of $2$ less than $2j+1,$ so $(x_{1}-x_{2})^{2^{a+1}}$ and $(x_{1}-x_{2})^{2^{a}}\prod(x_{i}-x_{j})^{2j+1-2^{a}}$ are both in $Q_{j}(3,\mathbf{F}_{2}).$ Further, $(x_{1}-x_{2})^{2^{a+1}}$ must be a generator and if $(x_{1}-x_{2})^{2^{a}}\prod(x_{i}-x_{j})^{2j+1-2^{a}}$ is a not a generator, by Lemma 3.4, $(x_{1}-x_{2})^{2^{a}}\prod(x_{i}-x_{j})^{2j+1-2^{a}}$ is generated by $(x_{1}-x_{2})^{2^{a+1}}$ which implies a relation between $(x_{1}-x_{2})^{2^{a+1}}$ and $(x_{1}-x_{2})^{2^{a}}\prod(x_{i}-x_{j})^{2j-2^{a}}.$ Because they freely generate $Q_{j-\frac{1}{2}}(3,\mathbf{F}_{2}),$ this is impossible. Thus $(x_{1}-x_{2})^{2^{a+1}}$ and $(x_{1}-x_{2})^{2^{a}}\prod(x_{i}-x_{j})^{2j+1-2^{a}}$ freely generate $Q_{m}(3,\mathbf{F}_{2})_{\mathrm{std}}$ by Lemma 3.5.
If $2j+1=2^{a+1},$ then both $(x_{2}-x_{3})^{2^{a+1}}$ and $(x_{2}-x_{3})^{2^{a+2}}$ lie in $Q_{j}(3,\mathbf{F}_{2}).$ The former is a generator by our inductive hypothesis. Since $2^{a+1}+2^{a+2}=6j+3,$ if the latter is not a generator, then by Lemma 3.4, $(x_{2}-x_{3})^{2^{a+2}}$ is generated by $(x_{2}-x_{3})^{2^{a+1}}$, which is false. Thus $(x_{2}-x_{3})^{2^{a+1}}$ and $(x_{2}-x_{3})^{2^{a+2}}$ freely generate $Q_{j}(3,\mathbf{F}_{2})$ by Lemma 3.5 as desired.
4 Properties of $3$ variable quasi-invariants
Similar to in the $p=2$ case, we can adapt many of the properties of $Q_{m}(3,\mathbf{F}_{p})$ for $p>3$ found in [wang2022explicit] to the $p=3$ case. We accomplish this by converting $\mathrm{std}$ to $\mathrm{sign-triv}$. For example, in $Q_{0}(3,\mathbf{F}_{3})$, the space spanned by $x_{1}-x_{2},x_{1}-x_{3}$ is a copy of $\mathrm{std}$. However, in $Q_{0}(3,\mathbf{F}_{3})$, the space spanned by $x_{1}-x_{2},x_{1}-x_{3}$ becomes a copy of $\mathrm{sign-triv}$. Using this we may show that there are equivalents of Lemma 3.2 to Lemma 3.5 from [wang2022explicit] in characteristic $3$.
We define $V_{ij}^{-}$ to be the $(-1)$-eigenspace of $s_{ij}$ in $V$, where $V$ is a copy of $\mathrm{std}$ or $\mathrm{sign-triv}$. Note that if $v\in V_{ij}^{-}$ we have $v=s_{23}v+s_{13}v$. The following lemma and corollary correspond to Lemma 3.2 and Corollary 3.3 respectively from [wang2022explicit].
Lemma 4.1.
[…]
$(x_{1}-x_{2})^{2m+1}K^{\prime}+(x_{2}-x_{3})^{2m+1}sK^{\prime}+(x_{3}-x_{1})^{2m+1}s^{2}K^{\prime}=0.$
Then $(x_{1}-x_{2})^{2m+1}K^{\prime}$ either belongs to $Q_{m}(3,\mathbf{F}_{3})_{\mathrm{sign}}$ or the $(-1)$-eigenspace of $s_{12}$ in some copy of $\mathrm{sign-triv}$ inside $Q_{m}(3,\mathbf{F}_{3})_{\mathrm{sign-triv}}$.
Proof.
The proof is largely the same as in [wang2022explicit]; the only difference is in the last step. Namely, now we have 2 2-dimensional indecomposable representations $\mathrm{sign-triv}$ and $\mathrm{triv-sign},$ but an element in the $(-1)$-eigenspace of $s_{12}$ in $\mathrm{triv-sign}$ must be in a copy of $\mathrm{sign}$. ∎
Corollary 4.2.
Let $K$ be a generator of $Q_{m}(3,\mathbf{F}_{3})$ and write $K=(x_{i}-x_{j})^{2m+1}K^{\prime}$. Then $K^{\prime}$ is not divisible by any nonconstant symmetric polynomial.
The proof of this corollary is identical to the proof of Corollary 3.3 in [wang2022explicit].
We define generators of $Q_{m}(3,\mathbf{F}_{3})$ to be “distinct” if they are either in different degrees, or if no linear combination of them over $\mathbf{F}_{3}$ is generated by lower degree generators.
Lemma 4.3.
Let $K$ and $L$ be distinct generators of $Q_{m}(3,\mathbbm{k})_{\mathrm{sign-triv}},$ and let $V$ and $W$ be the copies of $\mathrm{sign-triv}$ generated by $K$ and $L$ respectively. Then $Ks_{23}L+Ls_{23}K$ is a nonzero element of $Q_{m}(3,\mathbf{F}_{3})_{\mathrm{sign}}$ and $\deg V+\deg W\geq 6m+3$.
Noting that $\wedge^{2}(\mathrm{sign-triv})=\mathrm{sign}$, the proof of this lemma is also identical to the proof of Lemma 3.4 in [wang2022explicit].
Lemma 3.5 from [wang2022explicit] does not completely hold in characteristic $3$. A very similar and useful version does, however:
Lemma 4.4.
Assume that there exists generators $K$ and $L$ of $Q_{m}(3,\mathbf{F}_{3})_{\mathrm{sign-triv}}$ such that $\deg K+\deg L=6m+3$. Then $Q_{m}(3,\mathbf{F}_{3})_{\mathrm{sign-triv}}$ is freely generated by $K,$ $L,$ and $1$ over $\mathbf{F}_{3}[x_{1},x_{2},x_{3}]^{S_{3}}$.
Proof.
We note that $(L+s_{23}L)K-(K+s_{23}K)L=c\prod_{i<j}(x_{i}-x_{j})^{2m+1}$ for some $c\neq 0$ by Lemma 4.3. Moreover, $L+s_{23}L$ and $K+s_{23}K$ are symmetric because $K$ and $L$ are both acted on by negation by $s_{12},$ so elements in $Q_{m}(3,\mathbf{F}_{3})_{\mathrm{sign}}$ are generated by $K$ and $L.$ From there, the fact that $Q_{m}(3,\mathbf{F}_{3})_{\mathrm{sign-triv}}$ is generated by $K,$ $L,$ and $1$ over $\mathbf{F}_{3}[x_{1},x_{2},x_{3}]^{S_{3}}$ follows from the first part of the proof from [wang2022explicit].
To prove freeness, assume for the sake of contradiction that there was a relation $PK+QL+S=0$ for symmetric polynomials $P,$ $Q,$ and $S.$ $PK$ and $QL$ are both in the $(-1)$-eigenspace of $s_{12}$ while $S$ is not, so $S=0.$ Thus we have $PK=-QL$ and from there we can proceed the same as [wang2022explicit].
5 Ren–Xu counterexamples
We aim to explicitly describe the Hilbert series of $Q_{m}(3,\mathbf{F}_{3}).$ To do so we wish to identify the generators of $Q_{m}(3,\mathbf{F}_{3})_{\mathrm{sign-triv}}.$
In [Ren_2020], Ren and Xu found polynomials of the form $P_{k}^{3^{a}}\prod(x_{i}-x_{j})^{2b}$ in $Q_{m}(3,\mathbf{F}_{3})$ with degree less than $13m+$ 1 where $P_{k}$ is the map of the $13k+$ 1 degree generator of $Q_{k}(3,\mathbf{Q})$ into characteristic $3$ and where $a,$ $k,$ and $b$ are natural numbers. We refer to these polynomials as Ren–Xu counterexamples as they demonstrate the Hilbert series of $Q_{m}(3,\mathbf{F}_{3})$ differs from that of $Q_{m}(3,\mathbf{Q})$ for certain $m.$
Definition 5.1.
Let $\overline{P_{k}}$ be the generator of $Q_{k}(3,\mathbf{Q})$ of degree $13k+$ 1 in the $(-1)$-eigenspace of $s_{12}$, expressed as an integer polynomial with coprime coefficients. Let $P_{k}$ be the image of $\overline{P_{k}}$ under the quotient map $\mathbb{Z}[x_{1},x_{2},x_{3}]\to\mathbf{F}_{3}[x_{1},x_{2},x_{3}]$. Define the set $X$ as the set of all natural numbers $m$ such that $Q_{m}(3,\mathbf{F}_{3})$ has a Ren–Xu counterexample. Let $R_{m}$ be a lowest degree Ren–Xu counterexample in $Q_{m}(3,\mathbf{F}_{3})$ for all $m\in X.$
A key step in describing the Hilbert series of $Q_{m}(3,\mathbf{F}_{3})$ is proving [Ren_2020]’s conjecture:
Conjecture 5.2 ([Ren_2020]).
If the Hilbert series of $Q_{m}(3,\mathbf{F}_{p})$ differs from that of $Q_{m}(3,\mathbf{Q}),$ then there exists integers $a\geq 0$ and $k\geq 0$ such that
$\frac{mn(n-2)+\binom{n}{2}}{n(n-2)k+\binom{n}{2}-1}\leq p^{a}\leq\frac{mn}{nk+1}.$
The main step for proving the conjecture is the following theorem.
Theorem 5.3.
$Q_{m}(3,\mathbf{F}_{3})_{\mathrm{sign-triv}}$ is either freely generated by a generator of degree $3m+1,$ $23m+$ 2, and the polynomial $1$, or it is freely generated by $R_{m},$ another generator in degree $6m+3-\deg R_{m},$ and the polynomial $1$.
To prove this theorem, we first describe the Ren–Xu counterexamples.
Lemma 5.4.
If $m\in X,$ we must have $R_{m}=P_{k}^{3^{a}}\prod(x_{i}-x_{j})^{2b}$ where $a,b,k$ are natural numbers and $k\not\in X$.
Proof.
Assume for contradiction that there exists a nonnegative integer $m\in X$ such that $R_{m}=P_{k}^{3^{a}}\prod(x_{i}-x_{j})^{2b}$ where $a,b,k$ are natural numbers and $k\in X$. Then if $R_{k}=P_{l}^{3^{c}}\prod(x_{i}-x_{j})^{2d},$ the polynomial
${R_{k}}^{3^{a}}\prod(x_{i}-x_{j})^{2b}=P_{l}^{3^{a+c}}\prod(x_{i}-x_{j})^{2d\cdot 3^{a}+2b}$
has a strictly smaller degree than $R_{m}$ since $\deg R_{k}<3k+1=\deg P_{k}.$ Moreover, it is at least $m$-quasi-invariant, so it is a Ren–Xu counterexample for $Q_{m}(3,\mathbf{F}_{3})$. Yet $R_{m}$ is a minimal counterexample, giving a contradiction. ∎
This lemma allows us to consider only counterexamples $P_{k}^{3^{a}}\prod(x_{i}-x_{j})^{2b}$ such that $Q_{k}(3,\mathbf{F}_{3})$ does not contain a Ren–Xu counterexample.
From [Ren_2020], the Hilbert series for $Q_{m}(3,\mathbf{F}_{3})$ differs from characteristic 00 when there exists $a\in\mathbf{N}_{0}$ such that
$\frac{1}{3}\leq\left\{\frac{m}{3^{a}}\right\}\leq\frac{1}{3}-\frac{1}{3^{a}}.$
Notice this is equivalent to $m\pmod{3^{a}}$ being in $\{3^{a-1},3^{a-1}+1,\dotsm,2\cdot 3^{a-1}-1\}$.
Lemma 5.5.
If $m\not\in X$, then the base $3$ representation of $m$ contains no $1$’s.
Proof.
Suppose $m$ had the digit $1$ in the $a$th position from the right. Then $m\pmod{3^{a}}$ has a leading digit of $1$ if we choose $m\pmod{3^{a}}$ to be between 00 and $13^{a}-$ 1 inclusive. However, this implies that $m$ is in $\{3^{a-1},3^{a-1}+1,\dotsm,2\cdot 3^{a-1}-1\}$, so $m$ is a counterexample. ∎
Corollary 5.6.
If $m\not\in X$, then $m$ is even.
Proof.
From Lemma 5.5 $m$ has no $1$’s in its base $3$ representation, so
$m=\sum_{i=0}c_{i}3^{i}$
where $c_{i}$ is 00 or $2$. Thus $m$ must be even. ∎
Corollary 5.7.
For all $m\not\in X$, we have $m+1\in X$.
Proof.
By Corollary 5.6, if $m\not\in X,$ $m$ is even. Then $m+1$ is odd, so by the contrapositive of Corollary 5.6, $m+1\in X.$
Now we begin describing the degrees of Ren–Xu counterexamples.
Lemma 5.8.
If $Q_{m}(3,\mathbf{F}_{3})_{\mathrm{sign-triv}}$ has a generator in degree $3m+1,$ then $m+1\in X$ and $\deg R_{m+1}=3m+3$.
Proof.
If $m\in X,$ we must have $\deg R_{m}<3m+1.$ This implies a generator in a degree less than $13m+$ 1, violating Lemma 4.3. Thus $m\not\in X$, implying that $m+1\in X$ by Corollary 5.7.
Because $\deg R_{m+1}<3m+4$ and $Q_{m+1}(3,\mathbf{F}_{3})_{\mathrm{sign-triv}}\subset Q_{m}(3,\mathbf{F}_{3})_{\mathrm{sign-triv}}$, we have $43m+1\leq\deg R_{m+1}<3m+$ 4 . By construction $3|\deg R_{m+1},$ so $\deg R_{m+1}=3m+3.$
We now introduce a few useful lemmas.
Lemma 5.9.
Suppose $Q_{m}(3,\mathbf{F}_{3})_{\mathrm{sign-triv}}$ has a smallest degree generator $L$ in degree $13m+$ 1 . Assume that for all $i<m$, if $i\not\in X$, then $Q_{i}(3,\mathbf{F}_{3})_{\mathrm{sign-triv}}$ has a degree $13i+$ 1 generator. Then $Q_{m+1}(3,\mathbf{F}_{3})_{\mathrm{sign-triv}}$ has no nonsymmetric degree $13m+$ 1 or $23m+$ 2 element.
Proof.
Any nonsymmetric $13m+$ 1 degree element in $Q_{m+1}(3,\mathbf{F}_{3})_{\mathrm{sign-triv}}$ must be a scalar multiple of $L,$ so assume for contradiction $L$ is in $Q_{m+1}(3,\mathbf{F}_{3}).$ By Lemma 5.8, $R_{m+1}=P_{k}^{3^{a}}\prod(x_{i}-x_{j})^{2b}$ is in degree $33m+$ 3 for natural numbers $a,$ $b,$ $k.$ By Lemma 5.4, $k\not\in X$ implying $P_{k}$ is a $13k+$ 1 generator of $Q_{k}(3,\mathbf{F}_{3})_{\mathrm{sign-triv}}$ using our assumption. Moreover, with any other generator in a degree less than $33m+$ 3 violating Lemma 4.3, $R_{m+1}$ must be generated by $L,$ so $P_{k}^{3^{a}}\prod(x_{i}-x_{j})^{2b}=SL$ for some degree $2$ symmetric polynomial $S$. A degree $2$ symmetric polynomial divisible by $(x_{i}-x_{j})$ is impossible, so $S|P_{k}^{3^{a}}$ which implies either $S|P_{k}$ or $(x_{1}+x_{2}+x_{3})|P_{k}.$ If we let $P_{k}=P_{k}^{\prime}(x_{1}-x_{2})^{2k+1},$ in both cases $S|P_{k}^{\prime}$ or $(x_{1}+x_{2}+x_{3})|P_{k}^{\prime}$. However, by our assumption $P_{k}$ is a generator, so $P_{k}^{\prime}$ is not divisible by any nonconstant symmetric polynomial by Corollary 4.2.
Similarly, suppose for contradiction that $K$ is a nonsymmetric element of $Q_{m+1}(3,\mathbf{F}_{3})_{\mathrm{sign-triv}}$ of degree $23m+$ 2. Since $Q_{m+1}(3,\mathbf{F}_{3})_{\mathrm{sign-triv}}$ has no nonsymmetric $13m+$ 1 degree element, $K$ must be a generator. By Lemma 4.3, $K$ is the only generator in degree less than $3m+3,$ so $P_{k}^{3^{a}}\prod(x_{i}-x_{j})^{2b}$ is symmetric polynomial multiple of $K.$ However the only symmetric polynomials of degree $1$ are multiples of $x_{1}+x_{2}+x_{3},$ implying $(x_{1}+x_{2}+x_{3})|P_{k}$ which is impossible by Corollary 4.2. ∎
Note that by [felder2001action], $Q_{m}(3,\mathbf{Q})_{\mathrm{std}}$ has generators in degree $13m+$ 1 and $3m+2,$ and by [wang2022explicit], such generators with even degree are divisible by $x_{1}+x_{2}-2x_{3}.$ Let $\pi$ be the canonical mapping from characteristic 00 to characteristic $3$. We then have the following lemma:
Lemma 5.10.
Suppose $Q_{m}(3,\mathbf{F}_{3})_{\mathrm{sign-triv}}$ has a generator $L$ in degree $13m+$ 1 . We can choose the generators of $Q_{m}(3,\mathbf{Q})_{\mathrm{std}}$ to be integer polynomials $L^{\prime}$ and $(x_{1}+x_{2}-2x_{3})K^{\prime}$ with $\pi(K^{\prime})=\pi(L^{\prime})=L$. Moreover, if
$G=(x_{1}+x_{2}+x_{3})\left(\frac{K^{\prime}-L^{\prime}}{3}\right)-x_{3}K^{\prime},$
then
$\pi(G)=(x_{1}+x_{2}+x_{3})\pi\left(\frac{K^{\prime}-L^{\prime}}{3}\right)-x_{3}L$
is a degree $23m+$ 2 generator for $Q_{m}(3,\mathbf{F}_{3})_{\mathrm{sign-triv}}.$
Proof.
Let $L^{\prime}$ be an arbitrary $13m+$ 1 degree generator of $Q_{m}(3,\mathbf{Q})_{\mathrm{std}}$ with coprime integer coefficients in the $(-1)$-eigenspace of $s_{12}$. By Lemma 4.1, $\pi(L^{\prime})$ is an element of the $(-1)$-eigenspace of $s_{12}$ in $Q_{m}(3,\mathbf{F}_{3})_{\mathrm{sign-triv}}$ and if $\pi(L^{\prime})$ is not a scalar multiple of $L$ then there must exist some other generator of $Q_{m}(3,\mathbf{F}_{3})_{\mathrm{sign-triv}}$ with degree less than or equal to $13m+$ 1 . That generator and $L$ would violate Lemma 4.3, so we may set $\pi(L^{\prime})=L$.
A higher degree generator of $Q_{m}(3,\mathbf{Q})_{\mathrm{std}}$ has degree $23m+$ 2 . With $\deg L=3m+1$ implying $m\not\in X$, $23m+$ 2 is even by Corollary 5.6. Using [wang2022explicit], we let $(x_{1}+x_{2}-2x_{3})K^{\prime}$ be an arbitrary degree $23m+$ 2 generator for $Q_{m}(3,\mathbf{Q})_{\mathrm{std}}$ with coprime integer coefficients. Similarly $\pi((x_{1}+x_{2}-2x_{3})K^{\prime})=(x_{1}+x_{2}+x_{3})\pi(K^{\prime})$ is an element of $Q_{m}(3,\mathbf{F}_{3})_{\mathrm{sign-triv}},$ so $\pi(K^{\prime})$ is a non-symmetric polynomial of degree $13m+$ 1 in $Q_{m}(3,\mathbf{F}_{3})_{\mathrm{sign-triv}}$. Thus it must be a scalar multiple of $L,$ and we may set $\pi(K^{\prime})=L.$
Let $G=(x_{1}+x_{2}+x_{3})\left(\frac{K^{\prime}-L^{\prime}}{3}\right)-x_{3}K^{\prime}.$ Since
$(x_{1}+x_{2}-2x_{3})K^{\prime}-(x_{1}+x_{2}+x_{3})L^{\prime}=(x_{1}+x_{2}+x_{3})(K^{\prime}-L^{\prime})-3x_{3}K^{\prime}$
and $\pi(K^{\prime}-L^{\prime})=L-L=0,$ we have $G\in Q_{m}(3,\mathbf{Q})\cap\mathbf{Z}[x_{1},x_{2},x_{3}].$ Then
$\pi(G)=(x_{1}+x_{2}+x_{3})\pi\left(\frac{K^{\prime}-L^{\prime}}{3}\right)-x_{3}L.$
If $\pi(G)$ generated by $L$, we must have $\pi(G)=c(x_{1}+x_{2}+x_{3})L$ for some $c\in\mathbf{F}_{3}$ since $\deg\left(\pi(G)\right)=\deg(L)+1.$ However $x_{1}+x_{2}+x_{3}$ does not divide $x_{3}L$ since $L$ is a generator, so $x_{1}+x_{2}+x_{3}\nmid\pi(G).$ Then if $\pi(G)$ was not a generator, there must be some generator other than $L$ for $Q_{m}(3,\mathbf{F}_{3})$ in degree less than $23m+$ 2 which violates Lemma 4.3. Thus, $\pi(G)$ is a generator. ∎
We aim to prove that minimum Ren–Xu counterexamples are generators and represent the only cases where the Hilbert series of the quasi-invariants differs between characteristics 00 and $3$. To this end, we describe the degree of Ren–Xu counterexamples.
Example 5.11.
We notice a “staircase” pattern for Ren–Xu counterexamples. The following are counterexamples for $m=3,4,5:$
$(x_{1}-x_{2})^{9}$
$(x_{1}-x_{2})^{9}$
$(x_{1}-x_{2})^{9}\prod(x_{i}-x_{j})^{2}.$
We note that since $(x_{1}-x_{2})^{9}\in Q_{4}(3,\mathbf{F}_{3}),$ $(x_{1}-x_{2})^{9}$ is the Ren–Xu counterexample for both $m=3$ and $m=4.$ Moreover, the counterexample in $Q_{5}(3,\mathbf{F}_{3})$ is the previous counterexample $(x_{1}-x_{2})^{9}$ muliplied by $\prod(x_{i}-x_{j})^{2}$ to add the extra factor of $(x_{1}-x_{2})^{2}.$ In this way the degree of counterexample stays constant for the first half of the “staircase” and climbs by $6$ per each increase in $m$ thereafter. Moreover, we note that $m=2,6\not\in X,$ so our “staircase” is surrounded by non-counterexamples. One can also compute another generator for $m=3,$ $4,$ $5$ in degree $12,$ $18,$ and $18$ respectively. Since $9+12=6\cdot 3+3,$ $9+18=6\cdot 4+3,$ and $15+18=6\cdot 5+3,$ $Q_{m}(3,\mathbf{F}_{3})_{\mathrm{sign-triv}}$ is freely generated by each of these generators and $1$ by Lemma 4.4. This way we see that the upper degree generators form a complement to the lower degree ones, climbing by $6$ degrees initially and staying constant for the second half of the staircase.
Visually, the following figure shows the degree of the generators for $Q_{m}(3,\mathbf{F}_{3})$ with respect to $m$ were the staircase pattern and Theorem 5.3 to hold.
0010101010202020203030303040404040505050500050505050100100100100150150150150mDegree of generatorGenerator $1$Generator $2$
Figure 1: Degrees of generators in characteristic $3$ with respect to $m$
We prove that Ren–Xu counterexamples follow this staircase pattern:
Lemma 5.12.
Let $m$ be a natural number not in $X$ and let $d$ be the largest integer such that $R_{m+1}$ lies in $Q_{m+d}(3,\mathbf{F}_{3})$. Suppose that for all $k\leq m,$ if $k\not\in X$, then $Q_{k}(3,\mathbf{F}_{3})_{\mathrm{sign-triv}}$ has a generator in degree $13k+$ 1 . Then $R_{m+i}=R_{m+1}$ in degree $33m+$ 3 for $1\leq i\leq d$ and $R_{m+i}=R_{m+1}\prod(x_{i}-x_{j})^{2(i-d)}$ in degree $3m+3+6(i-d)$ for $d<i<2d.$
Proof.
Let $R_{m+1}=P_{k}^{3^{a}}\prod(x_{i}-x_{j})^{2b}$ where $k\in\mathbb{Z}_{\geq 0},a\in\mathbb{Z}_{>0}$, and $b=\max\left\{\frac{2m+3-3^{a}(2k+1)}{2},0\right\}.$ If $b$ is positive the polynomial $P_{k}^{3^{a}}\prod(x_{i}-x_{j})^{2(b-1)}$ has degree less than $23m-$ 2 and is at least $m$-quasi-invariant since $P_{k}^{3^{a}}\prod(x_{i}-x_{j})^{2b}$ has degree less than $43m+$ 4 . Thus $P_{k}^{3^{a}}\prod(x_{i}-x_{j})^{2(b-1)}$ is a Ren–Xu counterexample for $Q_{m}(3,\mathbf{F}_{3}),$ a contradiction.
In this way, we have $R_{m+1}=P_{k}^{3^{a}}.$ Moreover $Q_{k}(3,\mathbf{F}_{3})$ must be a non-counterexample by Lemma 5.4, so by our assumption $P_{k}$ is a generator. By Lemma 5.9, $P_{k}$ is not in $Q_{k+1}(3,\mathbf{F}_{3}),$ so the largest power of $(x_{1}-x_{2})$ dividing into $R_{m+1}$ must be $(x_{1}-x_{2})^{3^{a}(2k+1)}$ and $m+d=\frac{3^{a}(2k+1)-1}{2}$ by Lemma 4.1. Then for all $1\leq i\leq d,$
$\frac{2(m+i)+1-3^{a}(2k+1)}{2}$ $\leq\frac{2(m+d)+1-3^{a}(2k+1)}{2}$
$=0.$
Thus $R_{m+i}=P_{k}^{3^{a}}=R_{m+1}$ which is indeed in degree $33m+$ 3 by Lemma 5.8.
We claim that for $d<i<2d$, $m+i\in X$. Let $I$ be the set of integers $h$ such that $P_{k}^{3^{a}}\prod(x_{i}-x_{j})^{2b}$ is a Ren–Xu counterexample for $Q_{h}(3,\mathbf{F}_{3})$. Notice that if $o\in I,$ then the set $\{s,s+1,s+2,\ldots,s+3^{a-1}-1\}$ containing $o$ and for which $s\equiv 3^{a-1}\pmod{3^{a}}$ gives exactly $I.$ We then note that $m+1\in I,$ yet $m\not\in I$ since $m\not\in X$. Thus $m\equiv 3^{a-1}-1\pmod{3^{a}}.$ Since $m+d=\frac{3^{a}(2k+1)-1}{2}\in I$ as well, we have $s=3^{a}k+3^{a-1},$ $m=3^{a}k+3^{a-1}-1,$ and $d=\frac{3^{a-1}+1}{2}.$ Then
$\frac{3^{a}(2k+1)-1}{2}\leq m+i<\frac{3^{a-1}+1}{2}+\frac{3^{a}(2k+1)-1}{2}=3^{a}k+2\cdot 3^{a-1},$
so $m+i\pmod{3^{a}}$ is in $\{3^{a-1},3^{a-1}+1,\ldots,2\cdot 3^{a-1}-1\}$ and thus in $X$.
If $R_{m+i}=P_{k}^{3^{a}}\prod(x_{i}-x_{j})^{2b}$ where $b=\frac{2(m+i)+1-3^{a}(2k+1)}{2}$ for $d<i<2d,$ then $m+d=\frac{3^{a}(2k+1)-1}{2}$ implies $b=i-d.$ Thus $R_{m+i}=P_{k}^{3^{a}}\prod(x_{i}-x_{j})^{2(i-d)}$ has degree $3m+3+6(i-d)$ as desired.
In [wang2022explicit], the first author proved that generators $Q_{m}(3,\mathbf{F}_{p})_{\mathrm{std}}$ for $p>3$ lie in $\mathbf{F}_{p}[x_{1}-x_{3},x_{2}-x_{3}]$ using that $\mathbf{F}_{p}[x_{1}-x_{3},x_{2}-x_{3},x_{1}+x_{2}+x_{3}]=\mathbf{F}_{p}[x_{1},x_{2},x_{3}].$ However this is not true for $p=3$ since $x_{1}-x_{3}+x_{2}-x_{3}=x_{1}+x_{2}+x_{3}$ in characteristic $3,$ so we instead consider the space $\mathbf{F}_{3}[x_{1}-x_{3},x_{2}-x_{3},x_{3}].$ From now on we say a polynomial’s degree in $x_{3}$ is with respect to the basis $\{x_{1}-x_{3},x_{2}-x_{3},x_{3}\}.$ Moreover, in [wang2022explicit] the first author defined the polynomial $M_{d}=(x_{1}+x_{2}-2x_{3})^{2\left\{\frac{d}{2}\right\}}(x_{1}-x_{3})^{\bigl{\lfloor}\frac{d}{2}\bigr{\rfloor}}(x_{2}-x_{3})^{\bigl{\lfloor}\frac{d}{2}\bigr{\rfloor}}$ for natural numbers $d$ and proved that homogeneous $s_{12}$-invariant elements of $\mathbf{F}_{p}[x_{1}-x_{3},x_{2}-x_{3}]/(x_{1}-x_{2})^{2}$ are equal to constant multiples of $M_{d}.$ Extending this gives that elements of $\mathbf{F}_{3}[x_{1}-x_{3},x_{2}-x_{3},x_{3}]/(x_{1}-x_{2})^{2}$ are polynomials in $x_{3}$ with coefficients that are constant multiples of $M_{d}.$ This gives us intuition for following lemmas.
Lemma 5.13.
Let $e_{1},e_{2},$ and $e_{3}$ be the elementary symmetric polynomials for $\mathbf{F}_{3}[x_{1},x_{2},x_{3}]^{S_{3}}$ in degree $1,2,$ and $3$ respectively. If $n$ is a natural number such that $n\not\equiv 0\pmod{3},$ for all natural numbers $j<n$ there exists a monomial $P$ in $e_{1},e_{2},e_{3}$ such that $P$ has degree $n$ and degree $j$ in $x_{3}.$ If $n$ is a natural number such that $n\equiv 0\pmod{3},$ for all natural numbers $j<n-1$ there exists a monomial $P$ in $e_{1},e_{2},e_{3}$ such that $P$ has degree $n$ and degree $j$ in $x_{3}.$
Proof.
We choose $e_{1},$ $e_{2},$ and $e_{3}$ to be
$e_{1}=x_{1}+x_{2}+x_{3}=(x_{1}-x_{3})+(x_{2}-x_{3})$
$e_{2}=x_{1}x_{2}+x_{1}x_{3}+x_{2}x_{3}=(x_{1}-x_{3})(x_{2}-x_{3})+2((x_{1}-x_{3})+(x_{2}-x_{3}))x_{3}$
and
$e_{3}=x_{1}x_{2}x_{3}=(x_{1}-x_{3})(x_{2}-x_{3})x_{3}+((x_{1}-x_{3})+(x_{2}-x_{3}))x_{3}^{2}+x_{3}^{3}.$
We prove the lemma by induction on $j.$
The base case for $n$ where $3\nmid n$ is $j=n-1$. If $j=n-1$ and $n\equiv 1\pmod{3},$ we can let $P=e_{3}^{\frac{n-1}{3}}e_{1}.$ If $n\equiv 2\pmod{3},$ we let $P=e_{3}^{\frac{n-2}{3}}e_{2}.$ The base case when $3|n$ is $j=n-2,$ so we can let $P=e_{1}e_{2}e_{3}^{\frac{n}{3}-1}$
Suppose that, when $3\nmid n,$ for all $i$ such that $n>i>j$ where $j\in\mathbf{N}$ and $10\leq j<n-$ 1 there exists a monomial in $e_{1},e_{2},e_{3}$ with degree $n$ and degree $i$ in $x_{3}.$ Suppose the same for when $3|n$ but with $n-1>i>j$ and $j<n-2$. Then there exists a monomial $m=e_{1}^{a}e_{2}^{b}e_{3}^{c}$ with degree $j+1$ in $x_{3}$ in $\mathbf{F}_{3}[x_{1}-x_{3},x_{2}-x_{3},x_{3}]/(x_{1}-x_{2})^{2}.$ If $b\neq 0$ we can take the monomial $e_{1}^{a+2}e_{2}^{b-1}e_{3}^{c}$ to be $P$ since it has degree $n$ and degree $j$ in $x_{3}$. If $b=0$ and $a,c>0,$ then we take $P=e_{1}^{a-1}e_{2}^{b+2}e_{3}^{c-1}.$ Finally we are left with the cases $a,b=0$ or $b,c=0.$ The former would imply $m=e_{3}^{\frac{n}{3}}$ is our monomial, but $3\nmid n,$ would imply $m$ is not a polynomial and $3|n$ implies $m$ has degree $j+1=n$ in $x_{3}$ and $j=n-1\not<n-2.$ For the latter case, we have that $a=n,$ so $m=e_{1}^{n}$ implies that $j+1=0$ which is below our range for $j$.
Lemma 5.14.
For all $f_{i}\in\mathbf{F}_{3}$ and $n\not\equiv 0\pmod{3}$ there exists a $P\in\mathbf{F}_{3}[x_{1},x_{2},x_{3}]^{S_{3}}$ such that
$P=f_{0}M_{n}x_{3}^{0}+f_{1}M_{n-1}x_{3}^{1}+\dotsm+f_{n-2}M_{2}x_{3}^{n-2}+f_{n-1}M_{1}x_{3}^{n-1}$
in $\mathbf{F}_{3}[x_{1},x_{2},x_{3}]^{S_{3}}/(x_{1}-x_{2})^{2}.$ If $n\equiv 0\pmod{3},$ for all $f_{i}\in\mathbf{F}_{3}$ there exists a $P\in\mathbf{F}_{3}[x_{1},x_{2},x_{3}]^{S_{3}}$ such that
$P=f_{0}M_{n}x_{3}^{0}+f_{1}M_{n-1}x_{3}^{1}+\dotsm+f_{n-2}M_{2}x_{3}^{n-2}$
in $\mathbf{F}_{3}[x_{1},x_{2},x_{3}]^{S_{3}}/(x_{1}-x_{2})^{2}.$ Moreover, $P$ also satisfies the property that if it has degree $k$ in $x_{3}$ in $\mathbf{F}_{3}[x_{1},x_{2},x_{3}]^{S_{3}}/(x_{1}-x_{2})^{2},$ then it has degree $k$ in $x_{3}$ in $\mathbf{F}_{3}[x_{1}-x_{3},x_{2}-x_{3},x_{3}].$
Proof.
We prove this by induction.
For the base case when $n\not\equiv 0\pmod{3}$, we claim there exists coefficients $c_{i}\in\mathbf{F}_{3}$ such that the polynomial $c_{0}M_{n}x_{3}^{0}+c_{1}M_{n-1}x_{3}^{1}+\dotsm+c_{n-2}M_{2}x_{3}^{n-2}+c_{n-1}M_{1}x_{3}^{n-1}$ is in $\mathbf{F}_{3}[x_{1},x_{2},x_{3}]^{S_{3}}/(x_{1}-x_{2})^{2}.$ The symmetric polynomial 00 satisfies these conditions and has degree 00 in $x_{3}.$ For the base case when $n\equiv 0\pmod{3},$ we claim there exists coefficients $c_{0},...,c_{n-2}$ such that the polynomial $c_{0}M_{n}x_{3}^{0}+c_{1}M_{n-1}x_{3}^{1}+\dotsm+c_{n-2}M_{2}x_{3}^{n-2}$ is in $\mathbf{F}_{3}[x_{1},x_{2},x_{3}]^{S_{3}}/(x_{1}-x_{2})^{2}.$ The symmetric polynomial 00 satisfies this.
We consider the case where $n\not\equiv 0\pmod{3}.$ Suppose that for all $n\geq i>j$ there exists coefficients $c_{0},...,c_{i-1}$ such that for all $f_{i},f_{i+1},..,f_{n-1}$ there exists a symmetric polynomial $P$ such that
$P=c_{0}M_{n}x_{3}^{0}+c_{1}M_{n-1}x_{3}^{1}+\dotsm+c_{i-1}M_{n-i+1}x_{3}^{i-1}+f_{i}M_{n-i}x_{3}^{i}+\dotsm+f_{n-1}M_{1}x_{3}^{n-1}$
lies in $\mathbf{F}_{3}[x_{1},x_{2},x_{3}]^{S_{3}}/(x_{1}-x_{2})^{2}$ where $j\in\mathbf{N}$ and $10\leq j\leq n-$ 1 . Moreover, suppose the polynomial $P$ exists such that it has degree in $x_{3}$ equal to the degree in $x_{3}$ in $\mathbf{F}_{3}[x_{1},x_{2},x_{3}]^{S_{3}}/(x_{1}-x_{2})^{2}.$
Consider arbitrary coefficients $f_{j},f_{j+1},\ldots,f_{n-1}.$ If they are each $0,$ then we can take 00 to be our polynomial just like our base case. Otherwise, let $l$ be the greatest natural number $l\geq j$ such that $f_{l}\neq 0.$ If $l=j,$ by Lemma 5.13 there exists a monomial $m$ in $e_{1},e_{2},e_{3}$ with degree $j$ in $x_{3}$ and we may take $f_{j}m$ to be our symmetric polynomial.
If $l>j,$ by assumption there exists coefficients $c_{0},c_{1},\dots,c_{j}$ such that
$S=c_{0}M_{n}x_{3}^{0}+c_{1}M_{n-1}x_{3}^{1}+\dotsm+c_{j}M_{n-j}x_{3}^{j}+f_{j+1}M_{n-j-1}x_{3}^{j+1}+\dotsm+f_{n-1}M_{1}x_{3}^{n-1}$
lies in $\mathbf{F}_{3}[x_{1},x_{2},x_{3}]^{S_{3}}/(x_{1}-x_{2})^{2}.$ By assumption, $S$ has degree $l$ in $x_{3}.$
Without loss of generality let the leading coefficient of $m$ be $M_{n-j},$ so
$S+(f_{j}-c_{j})m=c_{0}^{\prime}M_{n}x_{3}^{0}+c_{1}^{\prime}M_{n-1}x_{3}^{1}+\dotsm+c_{j-1}^{\prime}M_{n-j+1}x_{3}^{j-1}+f_{j}M_{n-j}x_{3}^{j}+\dotsm+f_{n-1}M_{1}x_{3}^{n-1}$
for some coefficients $c_{0}^{\prime},c_{1}^{\prime},\ldots,c_{j-1}^{\prime}.$ Moreover $S+(f_{j}-c_{j})m$ is still a symmetric polynomial and $m$ has degree $j$ in $x_{3}$ while $S$ has degree $l,$ so $S+(f_{j}-c_{j})m$ has degree $l$ as desired.
An identical argument holds for $n\equiv 0\pmod{3}.$
Now we have the tools to prove $m\not\in X$ implies $m+1$ begins our staircase.
Lemma 5.15.
Suppose that for all $i\leq m,$ if $i\not\in X$ then $Q_{i}(3,\mathbf{F}_{3})$ has a $13i+$ 1 degree generator where $m$ is a natural number. Then if $Q_{m}(3,\mathbf{F}_{3})_{\mathrm{sign-triv}}$ has a generator in degree $13m+$ 1, $Q_{m+1}(3,\mathbf{F}_{3})_{\mathrm{sign-triv}}$ has a generator in degree $63m+$ 6 .
Proof.
By Lemma 5.10, the generators for $Q_{m}(3,\mathbf{F}_{3})_{\mathrm{sign-triv}}$ are
$\left((x_{1}+x_{2}+x_{3})\pi\left(\frac{A^{\prime}-B^{\prime}}{3}\right)-x_{3}B\right)(x_{1}-x_{2})^{2m+1}$
in degree $23m+$ 2, and
$B(x_{1}-x_{2})^{2m+1}$
in degree $13m+$ 1 where $(x_{1}-x_{2})^{2m+1}(x_{1}+x_{2}-2x_{3})A^{\prime}$ and $(x_{1}-x_{2})^{2m+1}B^{\prime}$ are the generators of $Q_{m}(3,\mathbf{Q})_{\mathrm{std}},$ $B$ is an $s_{12}$-invariant polynomial, and $\pi(A^{\prime})=\pi(B^{\prime})=B$.
For the greater degree generator, let $C=\left((x_{1}+x_{2}+x_{3})\pi\left(\frac{A^{\prime}-B^{\prime}}{3}\right)-x_{3}B\right)$. We would like to show there exists symmetric polynomials $P$ and $Q$ in degree $4$ and $5$ respectively such that
$PC+QB\equiv 0\pmod{(x_{1}-x_{2})^{2}}.$
This would then imply $(PC+QB)(x_{1}-x_{2})^{2m+1}\in Q_{m+1}(3,\mathbf{F}_{3})$ by Lemma 4.1. Consider writing
$P=f_{0}M_{4}x_{3}^{0}+f_{1}M_{3}x_{3}^{1}+f_{2}M_{2}x_{3}^{2}+f_{3}M_{1}x_{3}^{3}$
and
$Q=h_{0}M_{5}x_{3}^{0}+h_{1}M_{4}x_{3}^{1}+h_{2}M_{3}x_{3}^{2}+h_{3}M_{2}x_{3}^{3}+h_{4}M_{1}x_{3}^{4}$
for arbitrary $f_{i}$ and $h_{j}$. We know for any choice of $f_{i}$ and $h_{j},$ we have $P,Q\in\mathbf{F}_{3}[x_{1},x_{2},x_{3}]^{S_{3}}/(x_{1}-x_{2})^{2}$ by Lemma 5.14.
We claim that $B|\pi\left(\frac{A^{\prime}-B^{\prime}}{3}\right)$ in $\mathbf{F}_{3}[x_{1},x_{2},x_{3}]/(x_{1}-x_{2})^{2}.$ By [wang2022explicit] $A^{\prime}$ and $B^{\prime}$ are both polynomials in the variables $(x_{1}-x_{2})^{2}$ and $(x_{1}-x_{3})(x_{2}-x_{3})$. Moreover, by Lemma 5.9, $(x_{1}-x_{2})^{2}\nmid B$ so $B\equiv cM_{3m+1}\pmod{(x_{1}-x_{2})^{2}}$ for some $c\in\mathbf{F}_{3}$ such that $c\neq 0$. Similarly we know $\pi\left(\frac{A^{\prime}-B^{\prime}}{3}\right)\equiv c^{\prime}M_{3m+1}\pmod{(x_{1}-x_{2})^{2}}$ for some $c^{\prime}\in\mathbf{F}_{3}$. Thus we have $\pi\left(\frac{A^{\prime}-B^{\prime}}{3}\right)=dB$ where $d=\frac{c^{\prime}}{c}.$
Note that
$(x_{1}+x_{2}+x_{3})M_{2i}$ $=(x_{1}+x_{2}+x_{3})(x_{1}-x_{3})^{i}(x_{2}-x_{3})^{i}$
$=M_{2i+1}$
and
$(x_{1}+x_{2}+x_{3})M_{2i+1}$ $=(x_{1}+x_{2}+x_{3})^{2}(x_{1}-x_{3})^{i}(x_{2}-x_{3})^{i}$
$=((x_{1}-x_{2})^{2}+(x_{1}-x_{3})(x_{2}-x_{3}))(x_{1}-x_{3})^{i}(x_{2}-x_{3})^{i}$
$=(x_{1}-x_{3})(x_{2}-x_{3})M_{2i}$
$=M_{2i+2}.$
Using this we have
$PC+QB$ $=\left(h_{0}M_{5}B+f_{0}(x_{1}+x_{2}+x_{3})M_{4}\pi\left(\frac{A^{\prime}-B^{\prime}}{3}\right)x_{3}^{0}\right)$
$\quad+\sum_{i=1}^{3}\left(h_{i}M_{5-i}Bx_{3}^{i}+f_{i}(x_{1}+x_{2}+x_{3})M_{4-i}\pi\left(\frac{A^{\prime}-B^{\prime}}{3}\right)x_{3}^{i}-f_{i-1}M_{5-i}Bx_{3}^{i}\right)$
$\quad+h_{4}M_{1}Bx_{3}^{4}-f_{3}M_{1}Bx_{3}^{4}$
$=\left(h_{0}B+f_{0}\pi\left(\frac{A^{\prime}-B^{\prime}}{3}\right)\right)M_{5}$
$\quad+\sum_{i=1}^{3}\left(\left(\left(h_{i}-f_{i-1}\right)B+f_{i}\pi\left(\frac{A^{\prime}-B^{\prime}}{3}\right)\right)M_{5-i}x_{3}^{i}\right)+(h_{4}-f_{3})M_{1}Bx_{3}^{4}.$
$=\left(h_{0}+f_{0}d\right)BM_{5}+\sum_{i=1}^{3}\left(\left(h_{i}-f_{i-1}\right)+f_{i}d\right)BM_{5-i}x_{3}^{i}+(h_{4}-f_{3})M_{1}Bx_{3}^{4}.$
Letting $h_{i}$ be arbitrary for $i>0$, set $f_{3}=h_{4}$, $f_{i-1}=h_{i}+f_{i}d$ for $30<i<$ 3 and set $h_{0}=-f_{0}d$. This makes the expression $PC+QB=0.$
We claim $Q_{m+1}(3,\mathbf{F}_{3})_{\mathrm{sign-triv}}$ has a degree $33m+$ 3 generator, namely $R_{m+1}$. From Lemma 5.9, $Q_{m+1}(3,\mathbf{F}_{3})_{\mathrm{sign-triv}}$ has no degree $13m+$ 1 or $23m+$ 2 generator, so it has no generators in degree less than $33m+$ 3 . By Lemma 5.8, $R_{m+1}$ is in degree $33m+$ 3 so it must be a generator. Without loss of generality, we let
$R_{m+1}=\left((x_{1}+x_{2}+x_{3})C+SB\right)(x_{1}-x_{2})^{2m+1}$
where $S$ is a degree $2$ symmetric polynomial.
If $(PC+QB)(x_{1}-x_{2})^{2m+1}$ were generated by $R_{m+1},$ there exists a symmetric polynomial $I$ such that $IR_{m+1}=(PC+QB)(x_{1}-x_{2})^{2m+1}.$ This implies $(I(x_{1}+x_{2}+x_{3})-P)C+(IS-Q)B=0.$ If $I(x_{1}+x_{2}+x_{3})-P\neq 0$ or $IS-Q\neq 0,$ there is a relation on $C$ and $B$ over $\mathbf{F}_{3}[x_{1},x_{2},x_{3}]^{S_{3}}$, but $C(x_{1}-x_{2})^{2m+1}$ and $B(x_{1}-x_{2})^{2m+1}$ are generators of $Q_{m}(3,\mathbf{F}_{3}).$ Thus we must have $P=I(x_{1}+x_{2}+x_{3}),$ so $(x_{1}+x_{2}+x_{3})|P.$ Now we consider the symmetric polynomials $P^{\prime}=P+e_{2}^{2}+e_{2}e_{1}^{2}+e_{1}^{4}$ and $Q^{\prime}=Q+e_{3}e_{1}^{2}+(-d-1)e_{2}^{2}e_{1}-de_{1}^{3}e_{2}+(-d+1)e_{1}^{5}$. In $F_{3}[x_{1}-x_{3},x_{2}-x_{3},x_{3}]/(x_{1}-x_{2})^{2},$ we get that
$P^{\prime}=f_{0}M_{4}x_{3}^{0}+f_{1}M_{3}x_{3}^{1}+(f_{2}+1)M_{2}x_{3}^{2}+f_{3}M_{1}x_{3}^{3}$
and
$Q^{\prime}=h_{0}M_{5}x_{3}^{0}+h_{1}M_{4}x_{3}^{1}+(h_{2}-d)M_{3}x_{3}^{2}+(h_{3}+1)M_{2}x_{3}^{3}+h_{4}M_{1}x_{3}^{4}.$
Then $f_{2}+1=(h_{3}+f_{3}d)+1=(h_{3}+1)+f_{3}d,$ $f_{1}=h_{2}+f_{2}d=(h_{2}-d)+(f_{2}+1)d,$ and the rest of the equations necessary for $P^{\prime}C+Q^{\prime}B\equiv 0\pmod{(x_{1}-x_{2})^{2}}$ are the same as $PC+QB\equiv 0\pmod{(x_{1}-x_{2})^{2}}.$ Thus $P^{\prime}C+Q^{\prime}B\equiv 0\pmod{(x_{1}-x_{2})^{2}}.$ Moreover, $(x_{1}+x_{2}+x_{3})$ divides into $P+e_{2}e_{1}^{2}+e_{1}^{4}$ but not $e_{2}^{2},$ so $(x_{1}+x_{2}+x_{3})\nmid P^{\prime}.$ We have shown that if $(PC+QB)(x_{1}-x_{2})^{2m+1}$ is generated by $L,$ then $(x_{1}+x_{2}+x_{3})|P,$ implying $(P^{\prime}C+Q^{\prime}B)(x_{1}-x_{2})^{2m+1}$ is not generated by $R_{m+1}$. If $(P^{\prime}C+Q^{\prime}B)(x_{1}-x_{2})^{2m+1}$ is not a generator, then whatever generates it violates Lemma 4.3, so $(P^{\prime}C+Q^{\prime}B)(x_{1}-x_{2})^{2m+1}$ is indeed a degree $63m+$ 6 generator of $Q_{m+1}(3,\mathbf{F}_{3}).$
Now we prove that if $R_{m+1}$ begins our staircase, then it is the lower degree generator for the first half of the staircase.
Lemma 5.16.
Let $m\not\in X$ for some natural number $m.$ Suppose $R_{m+1}$ is a degree $33m+$ 3 generator of $Q_{m+1}(3,\mathbf{F}_{3})$ and $L$ is another generator in degree $63m+$ 6 . Further let $R_{m+1}$ lie in $Q_{m+d}(3,\mathbf{F}_{3})$ where $d$ is maximal. Then $L\prod(x_{i}-x_{j})^{2(i-1)},$ $R_{m+1},$ and $1$ freely generate $Q_{m+i}(3,\mathbf{F}_{3})_{\mathrm{sign-triv}}$ for $1\leq i\leq d.$
Proof.
$L\prod(x_{i}-x_{j})^{2(i-1)}$ lies in a copy of $\mathrm{sign-triv}$ and it is divisible by $(x_{1}-x_{2})^{2(m+i)+1},$ so it must lie in $Q_{m+i}(3,\mathbf{F}_{3})_{\mathrm{sign-triv}}$ with degree $3m+6i$ by Lemma 4.1. If $L\prod(x_{i}-x_{j})^{2(i-1)}$ is not a generator, $R_{m+1}$ must generate $L\prod(x_{i}-x_{j})^{2(i-1)},$ implying a relation between $R_{m+1}$ and $L.$ Thus $L\prod(x_{i}-x_{j})^{2(i-1)}$ is indeed a generator.
Moreover, $33m+3+3m+6i=6(m+i)+$ 3 so by Lemma 4.4, $L\prod(x_{i}-x_{j})^{2(i-1)}$ and $R_{m+1}$ generate $Q_{m+i}(3,\mathbf{F}_{3})_{\mathrm{sign-triv}}.$
Next, we prove that, for all consecutive spaces of quasi-invariants in the second half of the staircase, the lower degree generator is $\prod(x_{i}-x_{j})^{2}$ times the previous lower degree generator.
Lemma 5.17.
Let $m\not\in X$ for some natural number $m.$ Suppose $R_{m+1}$ is a degree $33m+$ 3 generator of $Q_{m}(3,\mathbf{F}_{3})$ and $L$ is another generator in degree $63m+$ 6 . Let $R_{m+1}$ lie in $Q_{m+d}(3,\mathbf{F}_{3})$ where $d$ is maximal. Further, let $L$ have degree $5$ in $x_{3}$. Then for all $d\leq i<2d,$ $Q_{m+i}(3,\mathbf{F}_{3})_{\mathrm{sign-triv}}$ is freely generated by a generator in degree $3m+6d,$ $R_{m+1}\prod(x_{i}-x_{j})^{2(i-d)}$ in degree $3m+6(i-d)+3,$ and $1$
Proof.
We proceed with induction.
The generator $R_{m+1}$ of $Q_{m+d}(3,\mathbf{F}_{3})$ is in degree $3m+3=3m+6(d-d)+3,$ and from Lemma 5.16 a second generator is $L\prod(x_{i}-x_{j})^{2(d-1)}$ in degree $3m+6d.$ Moreover these are the only generators so the claim is true for $i=d.$
Let $j$ be a natural number with $d<j<2d$ and suppose $Q_{m+i}(3,\mathbf{F}_{3})_{\mathrm{sign-triv}}$ has a generator in degree $3m+6d$ and degree $33m+6(i-d)+$ 3 for all $d\leq i<j$ where this upper degree generator is a polynomial of degree at most $5$ in $x_{3}$ and is not generated by $R_{m+1}$. Consider $Q_{m+j}(3,\mathbf{F}_{3})_{\mathrm{sign-triv}}.$ We know $R_{m+1}\prod(x_{i}-x_{j})^{2(j-d-1)}$ is an element of $Q_{m+j-1}(3,\mathbf{F}_{3})_{\mathrm{sign-triv}}$ of degree $33m+6(j-d-1)+$ 3 by Lemma 4.1. Since $j-1<j$ we may use our inductive hypothesis implies $R_{m+1}\prod(x_{i}-x_{j})^{2(j-d-1)}$ is a generator for $Q_{m+j-1}(3,\mathbf{F}_{3})_{\mathrm{sign-triv}}$.
Let $T$ be the degree $3m+6d$ generator for $Q_{m+j-1}(3,\mathbf{F}_{3})_{\mathrm{sign-triv}}$ with degree $5$ in $x_{3}.$ We write $R_{m+1}\prod(x_{i}-x_{j})^{2(j-d-1)}=R_{m+1}^{\prime}(x_{1}-x_{2})^{2(m+j-1)+1}$ and $T=T^{\prime}(x_{1}-x_{2})^{2(m+j-1)+1}$ for $s_{12}$ invariant polynomials $R_{m+1}^{\prime}$ and $T^{\prime}$. If $o=m+4j-6d-2$ and $r=m+6d-2j+1,$ then $\deg R_{m+1}^{\prime}=o$ and $\deg T^{\prime}=r.$ We want to find a degree $r-o$ symmetric polynomial $P$ such that
$-PR_{m+1}^{\prime}+T^{\prime}\equiv 0\pmod{(x_{1}-x_{2})^{2}}.$
We claim that $R_{m+1}^{\prime}$ is degree 00 in $x_{3}.$ This is because $R_{m+1}\prod(x_{i}-x_{j})^{2(j-d-1)}=P_{l}^{3^{a}}\prod(x_{i}-x_{j})^{2(j-d-1)}$ as we proved in Lemma 5.12. Since $P_{l}$ is the map of the generator of $Q_{l}(3,\mathbf{Q})$ into characteristic $3,$ $P_{l}$ must be constant in the variable $x_{3}.$ We can see $\prod(x_{i}-x_{j})^{2(j-d-1)}$ is also constant in $x_{3},$ so $R_{m+1}$ and $R_{m+1}^{\prime}$ are constant in $x_{3}.$
Having assumed that $T^{\prime}$ is at most degree $5$ in $x_{3},$
$T^{\prime}=t_{0}M_{r}x_{3}^{0}+t_{1}M_{r-1}x_{3}^{1}+t_{2}M_{r-2}x_{3}^{2}+t_{3}M_{r-3}x_{3}^{3}+t_{4}M_{r-4}x_{3}^{4}+t_{5}M_{r-5}x_{3}^{5}$
and
$R_{m+1}^{\prime}=aM_{o}$
for coefficients $t_{i}$ and $a$ in $\mathbf{F}_{3}.$ Since $R_{m+1}$ is not in $Q_{m+d+1}(3,\mathbf{F}_{3})_{\mathrm{sign-triv}},$ we have $a\neq 0.$ We let
$P=\frac{t_{0}}{a}M_{r-o}x_{3}^{0}+\frac{t_{1}}{a}M_{r-o-1}x_{3}^{1}+\frac{t_{2}}{a}M_{r-o-2}x_{3}^{2}+\frac{t_{3}}{a}M_{r-o-3}x_{3}^{3}+\frac{t_{4}}{a}M_{r-o-4}x_{3}^{4}+\frac{t_{5}}{a}M_{r-o-5}x_{3}^{5}$
so that $T^{\prime}-PR_{m+1}^{\prime}\equiv 0\pmod{(x_{1}-x_{2})^{2}}.$ Since $\deg(P)=r-o=12d-6j+3\geq 9>7$, by Lemma 5.14 such a symmetric polynomial $P$ is attainable with $P$ having degree at most degree $5$ in $x_{3}.$ Since $T^{\prime}$ also has at most degree $5$ in $x_{3}$ and $R_{m+1}^{\prime}$ has degree $0,$ $(-PR_{m+1}^{\prime}+T^{\prime})$ has at most degree $5$ in $x_{3}$. Letting $U=(-PR_{m+1}^{\prime}+T^{\prime})(x_{1}-x_{2})^{2(m+j-1)+1}$, we have $U$ is in $Q_{m+j}(3,\mathbf{F}_{3})$ with degree $3m+6d$ and since $(-PR_{m+1}^{\prime})(x_{1}-x_{2})^{2(m+j-1)+1}$ is generated by $R_{m+1}$ and $T$ is not, $U$ is not generated by $R_{m+1}$. Finally we also have $R_{m+1}\prod(x_{i}-x_{j})^{2(j-d)}$ is in $Q_{m+j}(3,\mathbf{F}_{3})_{\mathrm{sign-triv}}$ with degree $33m+6(j-d)+$ 3 . Thus what is left is to prove is $R_{m+1}\prod(x_{i}-x_{j})^{2(j-d)}$ and $(-PR_{m+1}^{\prime}+T^{\prime})(x_{1}-x_{2})^{2(m+j-1)+1}$ are generators for $Q_{m+j}(3,\mathbf{F}_{3}).$
Assume for sake of contradiction that $U$ and $R_{m+1}\prod(x_{i}-x_{j})^{2(j-d)}$ are not both generators. If $R_{m+1}\prod(x_{i}-x_{j})^{2(j-d)}$ is a generator, then any other generator must be of at least degree $3m+6d$ by Lemma 4.3. Yet $U$ is not generated by $R_{m+1}\prod(x_{i}-x_{j})^{2(j-d)}$ since it is not generated by $R_{m+1}.$ Thus $U$ must be a generator.
Next, we consider if $R_{m+1}\prod(x_{i}-x_{j})^{2(j-d)}$ is not a generator. For $R_{m+1}\prod(x_{i}-x_{j})^{2(j-d)}$ to not be a generator there must be a generator in a degree less than $33m+6(j-d)+$ 3. Let it be $G$, and by Lemma 4.3, any other generator must have degree greater than $3m+6d.$ Thus $U$ is not a generator, so $U$ and $R_{m+1}\prod(x_{i}-x_{j})^{2(j-d)}$ are both generated by $G$ and specifically $U=QG$ and $R_{m+1}\prod(x_{i}-x_{j})^{2(j-d)}=SG$ for symmetric polynomials $P$ and $Q.$ Moreover, $R_{m+1}\prod(x_{i}-x_{j})^{2(j-d-1)}$ is the lowest degree generator for $Q_{m+j-1}(3,\mathbf{F}_{3})_{\mathrm{sign-triv}},$ so $G=CR_{m+1}\prod(x_{i}-x_{j})^{2(j-d-1)}$ for a symmetric polynomial $C.$ This implies $C|\prod(x_{i}-x_{j})^{2},$ and $G$ is not a scalar multiple of $R_{m+1}\prod(x_{i}-x_{j})^{2(j-d)},$ so $C$ is a constant. We then have $U$ is a constant multiple of $QR_{m+1}\prod(x_{i}-x_{j})^{2(j-d-1)},$ so $U$ is generated by $R_{m+1}$ which is a contradiction.
Thus $U$ and $R_{m+1}\prod(x_{i}-x_{j})^{2(j-d)}$ are each generators and together with $1$ they freely generate $Q_{m}(3,\mathbf{F}_{3})_{\mathrm{sign-triv}}$ by Lemma 4.4
Finally, we show that after the staircase completes, the next space of quasi-invariants has no counterexamples.
Lemma 5.18.
Let $Q_{m-1}(3,\mathbf{F}_{3})_{\mathrm{sign-triv}}$ have generators $K$ in degree $33m-$ 3 and $T$ in degree $3m$ such that $K$ is not in $Q_{m}(3,\mathbf{F}_{3})_{\mathrm{sign-triv}}.$ If $m$ is even then $Q_{m}(3,\mathbf{F}_{3})_{\mathrm{sign-triv}}$ is freely generated by a generator in degree $3m+1,$ $3m+2,$ and $1$
Proof.
Suppose for the sake of contradiction that $Q_{m}(3,\mathbf{F}_{3})_{\mathrm{sign-triv}}$ has a generator in degree $3m.$ Without loss of generality let that generator be $T.$ From [felder2001action], we can let $L^{\prime}$ be a degree $13m+$ 1 generator of $Q_{m}(3,\mathbf{Q})_{\mathrm{std}}$ with coprime integer coefficients. Then $\pi(L^{\prime})\in Q_{m}(3,\mathbf{F}_{3})_{\mathrm{sign-triv}},$ so $\pi(L^{\prime})$ must be generated by $T$ since any other generator in degree less than degree $13m+$ 1 would violate Lemma 4.3. Moreover, the only degree $1$ symmetric polynomials are constant multiples of $x_{1}+x_{2}+x_{3},$ so we can assume without loss of generality that
$\pi(L^{\prime})=(x_{1}+x_{2}+x_{3})T.$
Note that from [wang2022explicit] all generators of $Q_{m}(3,\mathbf{Q})_{\mathrm{std}}$ must lie in $\mathbf{Q}[x_{1}-x_{3},x_{2}-x_{3}].$ Thus $(x_{1}+x_{2}+x_{3})T\in\mathbf{F}_{3}[x_{1}-x_{3},x_{2}-x_{3}]$ and so $T\in\mathbf{F}_{3}[x_{1}-x_{3},x_{2}-x_{3}].$
We also have $T=(x_{1}-x_{2})^{2m+1}T^{\prime}$ for some $s_{12}$-invariant polynomial $T^{\prime}$. Thus by the fundamental theorem of symmetric polynomials $T^{\prime}\in\mathbf{F}_{3}[(x_{1}-x_{3})(x_{2}-x_{3}),x_{1}+x_{2}+x_{3}].$ Note that $\deg T^{\prime}=3m-2m-1=m-1$ and $m$ is even, so $T^{\prime}$ has an odd degree. However, since it is generated by $(x_{1}-x_{3})(x_{2}-x_{3})$ and $x_{1}+x_{2}+x_{3},$ we must have $(x_{1}+x_{2}+x_{3})|T^{\prime}.$ This gives a contradiction because $T$ is a generator.
Finally, we have the lemmas to prove Theorem 5.3.
Proof of Theorem 5.3.
We prove this using induction on $m.$
The generators for $Q_{0}(3,\mathbf{F}_{3})_{\mathrm{sign-triv}}$ are $x_{1}-x_{2}$ and $x_{3}(x_{1}-x_{2}).$ These generators are in degree $3\cdot 0+1$ and $3\cdot 0+2$ so the theorem is true for the base case.
Assume the claim is true when $m<j$ for some $j\in\mathbf{N}.$ Consider the space $Q_{j}(3,\mathbf{F}_{3})_{\mathrm{sign-triv}}$. Let $t$ be the largest natural number less than $j$ such that $t\not\in X$. By the inductive hypothesis $Q_{t}(3,\mathbf{F}_{3})$ has a generator in degree $13t+$ 1 and $23t+$ 2 . By Lemma 5.10, we may let the generators be
$\left((x_{1}+x_{2}+x_{3})\pi\left(\frac{A^{\prime}-B^{\prime}}{3}\right)-x_{3}B\right)(x_{1}-x_{2})^{2m+1}$
and
$B(x_{1}-x_{2})^{2m+1}$
where $(x_{1}-x_{2})^{2m+1}(x_{1}+x_{2}-2x_{3})A^{\prime}$ and $(x_{1}-x_{2})^{2m+1}B^{\prime}$ are generators for $Q_{t}(3,\mathbf{Q})_{\mathrm{std}}$ and $\pi(A^{\prime})=\pi(B^{\prime})=B.$ From Lemma 5.15, $Q_{t+1}(3,\mathbf{F}_{3})_{\mathrm{sign-triv}}$ is generated by a generator in degree $63t+$ 6 and $33t+$ 3 . Moreover, $R_{t+1}$ is the $33t+$ 3 degree generator by Lemma 5.8. Let $L$ be the degree $63t+$ 6 generator. Suppose $R_{t+1}$ lies in $Q_{t+d}(3,\mathbf{F}_{3}),$ but not $Q_{t+d+1}(3,\mathbf{F}_{3})$ where $d$ is a natural number.
First we consider when $t+d\geq j\geq t+1.$ By Lemma 5.16, $Q_{j}(3,\mathbf{F}_{3})_{\mathrm{sign-triv}}$ has generators $R_{t+1}$ and $L\prod(x_{i}-x_{j})^{2(j-t-1)}.$ Note that $R_{t+1}=R_{j}$ by Lemma 5.12, and further $\deg(L)+\deg(R_{t+1})=(6(j-t-1)+3t+6)+3t+3=6j+3.$ By Lemma 4.4, we then have that $R_{t+1}$ and $L\prod(x_{i}-x_{j})^{2(j-t-1)}$ generate $Q_{j}(3,\mathbf{F}_{3})_{\mathrm{sign-triv}}.$
Next, we consider the case where $t+2d-1\geq j\geq t+d+1.$ Notice that by our construction in Lemma 5.15, we can choose $L$ such that it has at most degree $5$ in $x_{3}.$ Thus we can apply Lemma 5.17, which gives us that $Q_{j}(3,\mathbf{F}_{3})_{\mathrm{sign-triv}}$ is generated by $R_{t+1}\prod(x_{i}-x_{j})^{2(j-t-d)}$ and a generator in degree $3t+6d.$ Note that $R_{t+1}\prod(x_{i}-x_{j})^{2(j-t-d)}$ is a constant multiple of $R_{j}$ by Lemma 5.12. Moreover, the sum of their degrees is $33t+6(j-t-d)+3+3t+6d=6j+$ 3 as desired.
Finally we consider if $j=t+2d.$ Note that by Lemma 5.17, $Q_{t+2d-1}(3,\mathbf{F}_{3})$ has a generator in degree $3t+6d$ and $33t+6(d-1)+$ 3 . The degree $33t+6(d-1)+$ 3 generator is $R_{t+1}\prod(x_{i}-x_{j})^{2(d-1)},$ and $R_{t+1}$ is divisible by $(x_{1}-x_{2})^{2(t+d)+1}$ where $d$ is maximal, so $R_{t+1}\prod(x_{i}-x_{j})^{2(d-1)}$ does not lie in $Q_{t+2d}(3,\mathbf{F}_{3}).$ Moreover, $Q_{t}(3,\mathbf{F}_{3})$ is a non Ren–Xu counterexample, so $t$ must be even by Lemma 5.6. Then $t+2d$ is even as well, so by Lemma 5.18, $Q_{t+2d}(3,\mathbf{F}_{3})$ has a generator in degree $13(t+2d)+$ 1 and $23(t+2d)+$ 2 .
Now we claim we have exhausted all cases. If we had $j>t+2d,$ since we just showed $t+2d\not\in X,$ we would not have chosen $t$ to be the largest natural number less than $j$ not in $X.$
Remark 5.19.
We can compute the degrees of generators of $Q_{m}(3,\mathbf{F}_{3})_{\mathrm{sign-triv}}$ explicitly. If $m$ has no digits $1$ in its base $3$ representation, then the generators have degree $13m+$ 1 and $23m+$ 2 . Otherwise the lower degree generator is $R_{m}.$ We can deduce the minimal degree of the Ren–Xu counterexamples in $Q_{m}(3,\mathbf{F}_{3}):$ Let $a$ be the greatest natural number such that the $a$th term from the right in the base $3$ representation of $m$ is $1$ Then if $\left\lceil\frac{\left\lceil\frac{m}{3^{a}}\right\rceil-1}{2}\right\rceil=k,$ a minimal degree Ren–Xu counterexample is $P_{k}^{3^{a}}\prod(x_{i}-x_{j})^{2b}$ where $b=\max\left\{\frac{2m+1-3^{a}(2k+1)}{2},0\right\}.$ The degrees of the generators are then $3^{a}(2k+1)+6b$ and $6m+3-3^{a}(2k+1)-6b$.
6 Representations of $S_{3}$ in $Q_{m}(3,\mathbf{F}_{3})$
Now that we have a complete picture of $Q_{m}(3,\mathbf{F}_{3})_{\mathrm{sign-triv}}$, we consider generators that generate the other indecomposable modules of $S_{3}$. We start with $\mathrm{triv-sign-triv}$, which behaves very similarly to $\mathrm{sign-triv}$.
Proposition 6.1.
Suppose that for all $i\leq m$, $Q_{i}(3,\mathbf{F}_{3})_{\mathrm{sign-triv}}$ has generators in degree $d$ and $6i+3-d$ respectively for some $d$. If $K,L$ are distinct generators of $Q_{m}(3,\mathbf{F}_{3})_{\mathrm{sign-triv}}$ then there are two other homogeneous generators $K_{1},L_{1}$ of $Q_{m}(3,\mathbf{F}_{3})$ in the same degrees as $K,L$, respectively such that as a representation of $S_{3}$, $K_{1}$ generates a copy of $\mathrm{triv-sign-triv}$ containing $K$ and $L_{1}$ generates a copy of $\mathrm{triv-sign-triv}$ containing $L$. Moreover, there are no relations between $K_{1},L_{1}$ over the symmetric polynomials, and there are no other generators of $Q_{m}(3,\mathbf{F}_{3})$ that generate a copy of $\mathrm{triv-sign-triv}$.
Proof.
We prove this by induction on $m$. For the base case $m=0$, note that by Example 2.8, for $K=x_{1}-x_{2}$ we have that $K_{1}=x_{1}$ satisfies the desired conditions. Similarly, for $L=(x_{1}-x_{2})x_{3}$, we have that $L_{1}=x_{1}(x_{2}+x_{3})$ satisfies the desired conditions. These two are independent over the symmetric polynomials, as a relation between them would imply a relation between 1 and $x_{2}+x_{3}$.
For the inductive step, let $K^{\prime},L^{\prime}$ be the generators of $Q_{m-1}(3,\mathbf{F}_{3})_{\mathrm{sign-triv}}$ and let $K^{\prime}_{1},L^{\prime}_{1}$ be the corresponding generators of $Q_{m-1}(3,\mathbf{F}_{3})$. Without loss of generality, we can choose $K^{\prime}_{1},L^{\prime}_{1}$ to be $s_{23}$-invariant with $(1-s_{12})K^{\prime}_{i}=K^{\prime},(1-s_{12})L^{\prime}_{i}=L^{\prime}$ (similar to in the base case). Let $K,L$ be generators of $Q_{m}(3,\mathbf{F}_{3})_{\mathrm{sign-triv}}$. Then since $K,L\in Q_{m-1}(3,\mathbf{F}_{3})_{\mathrm{sign-triv}}$, we can write $K=P_{1}K^{\prime}+Q_{1}L^{\prime}$, $L=P_{2}K^{\prime}+Q_{2}L^{\prime}$ for symmetric polynomials $P_{1},P_{2},Q_{1},Q_{2}$. Then it follows that $K_{1}:=P_{1}K^{\prime}_{1}+Q_{1}L^{\prime}_{1},L_{1}:=P_{2}K^{\prime}_{1}+Q_{2}L^{\prime}_{1}$ each generate a copy of $\mathrm{triv-sign-triv}$ that contains $K,L$, respectively. Moreover, if there is some relation $P_{3}K_{1}+Q_{3}L_{1}=0$ for symmetric polynomials $P_{3},Q_{3}$, then applying $1-s_{12}$ to this equation would yield $P_{3}K+Q_{3}L=0$, which violates Lemma 4.4.
Next, we show that $K_{1},L_{1}$ are $m$-quasi-invariant. As the computations are the same for both polynomials, we give the proof only for $K_{1}$. First, note that $(1-s_{23})K_{1}=0$ since both $K^{\prime}_{1},L^{\prime}_{1}$ are $s_{23}$-invariant. Next, note that $(1-s_{12})K_{1}=K$ is divisible by $(x_{1}-x_{2})^{2m+1}$ by Lemma 4.1. Finally, note that since $K_{1}$ is $s_{23}$-invariant, we have
$(1-s_{13})K_{1}=s_{23}(s_{23}-s_{23}s_{13})K_{1}=s_{23}(1-s_{23}s_{13}s_{23})K_{1}=s_{23}(1-s_{12})K_{1}$
is divisible by $s_{23}(x_{1}-x_{2})^{2m+1}=(x_{1}-x_{3})^{2m+1}$.
Note that $K_{1},L_{1}$ are the minimal degree polynomials such that $(1-s_{12})K_{1},(1-s_{12})L_{1}$ are symmetric polynomial multiples of $K,L$, respectively, so they cannot be generated by any other generators and thus must be a generators themselves. Then assume for contradiction that there is some other generator $T$ of $Q_{m}(3,\mathbf{F}_{3})$ that generates a copy of $\mathrm{triv-sign-triv}$. Then $(1-s_{12})T$ is contained in a copy of $\mathrm{sign-triv}$ and is $s_{12}$-antiinvariant, so we can write $(1-s_{12})T=S_{1}K+S_{2}L$ for symmetric polynomials $S_{1},S_{2}$. Then $T,S_{1}K_{1}+S_{2}L_{1}$ generate copies of $\mathrm{triv-sign-triv}$ with the same $\mathrm{sign-triv}$ submodule, so they generate a copy of
$(\mathrm{triv-sign-triv}\oplus\mathrm{triv-sign-triv})/\mathrm{sign-triv}\cong\mathrm{triv-sign-triv}\oplus\mathrm{triv}.$
Thus $T$ is generated by $K_{1},L_{1},1$, and is not a generator itself. ∎
Corollary 6.2.
The generators $1,K,K_{1},L,L_{1}$ of $Q_{m}(3,\mathbf{F}_{3})$ defined in Proposition 2.9, Theorem 5.3, and Proposition 6.1 have no relations between them over the symmetric polynomials.
Proof.
Let
$P_{1}+P_{2}K+P_{3}L+P_{4}K_{1}+P_{5}L_{1}=0$
for symmetric polynomials $P_{1},\dots,P_{5}$. Then apply $1+s_{12}$ to the equation to yield
$02P_{1}+P_{4}(2K_{1}-K)+P_{5}(2L_{1}-L)=$ 0
since $K,L$ are $s_{12}$-antiinvariant. Next, apply $1-s_{23}$ to this equation to yield
$P_{4}(s_{23}-1)K+P_{5}(s_{23}-1)L=0.$
Note that $(s_{23}-1)K$ generates the same copy of $\mathrm{sign-triv}$ as $K$, since $s_{23}-1$ acts bijectively on $\mathrm{sign}$ (and similarly for $L$). So a relation between $(s_{23}-1)K,(s_{23}-1)L$ is equivalent to a relation between $K,L$, which cannot exist by Lemma 4.4. So we have $P_{4}=P_{5}=0$.
Now, the result follows from Lemma 4.4. ∎
Remark 6.3.
In the non-modular case, one has that the polynomial $\prod_{i<j}(x_{i}-x_{j})^{2m+1}$ is a generator of $Q_{m}(n,\mathbbm{k})$, as it is the lowest degree quasi-invariant in the sign module. However, from Lemma 4.4 we have that in characteristic 3,
$(L+s_{23}L)K-(K+s_{23}K)L=c\prod_{i<j}(x_{i}-x_{j})^{2m+1},$
so $\prod_{i<j}(x_{i}-x_{j})^{2m+1}$ is not a generator. We can take this calculation further, and note that $(L+s_{23}L)K_{1}-(K+s_{23}K)L_{2}$ would then generate a copy of $\mathrm{triv-sign}$, as the quotient of this module by the space generated by $(L+s_{23}L)K-(K+s_{23}K)L$ must be a trivial module.
Now, we now only need to consider the modules $\mathrm{triv-sign},\mathrm{sign-triv-sign}$. To motivate the results that follow, we start by considering 0-quasi-invariants.
Example 6.4.
Note that from Corollary 6.2 we know that $Q_{0}(3,\mathbf{F}_{3})$ has 5 generators $1,x_{1}-x_{2},(x_{1}-x_{2})x_{3},x_{1},x_{1}(x_{2}+x_{3})$ with no relations between them. By examining the dimension of the space of all homogeneous degree 3 polynomials, we have that $Q_{0}(3,\mathbf{F}_{3})[3]$ is 10-dimensional, and so far we have accounted for $3+2+2+1+1=9$ dimensions. Moreover, every irreducible representation is accounted for, so this extra dimension must be an extension of an existing indecomposable representation. The only indecomposable representations that have nontrivial extensions are the triv generated by $x_{1}x_{2}x_{3}$ and the $\mathrm{triv-sign}$ generated by
$E:=(x_{1}x_{2}+x_{1}x_{3}+x_{2}x_{3})x_{1}+(x_{1}+x_{2}+x_{3})(x_{1}(x_{2}+x_{3}))=-x_{1}^{2}x_{2}-x_{1}^{2}x_{3}+x_{1}x_{2}^{2}+x_{1}x_{3}^{2}.$
Indeed, the $\mathrm{triv-sign}$ generated by $E$ extends to a $\mathrm{sign-triv-sign}$ generated by
$F:=(x_{1}-x_{2})x_{1}x_{2}.$
We will later see that the polynomials $E,F$ defined above are key to understanding $\mathrm{triv-sign}$ and $\mathrm{sign-triv-sign}$ in the quasi-invariants.
Proposition 6.5.
$Q_{0}(3,\mathbf{F}_{3})$ is freely generated by $1,x_{1}-x_{2},(x_{1}-x_{2})x_{3},x_{1},x_{1}(x_{2}+x_{3}),F$ as a $\mathbf{F}_{3}[x_{1},x_{2},x_{3}]^{S_{3}}$-module.
Proof.
We already know that the first 5 polynomials are independent. Now, let
$P_{1}+P_{2}(x_{1}-x_{2})+P_{3}(x_{1}-x_{2})x_{3}+P_{4}x_{1}+P_{5}(x_{2}+x_{3})x_{1}+P_{6}F=0$
for symmetric polynomials $P_{i}$. Apply $1-s_{12}$ to this equation to get
$(P_{4}-P_{2})(x_{1}-x_{2})+(P_{5}-P_{3})(x_{1}-x_{2})x_{3}-P_{6}F=0.$
Next, apply $1+s_{23}$ to get
$(P_{2}-P_{4})(x_{1}+x_{2}+x_{3})+(P_{5}-P_{3})(x_{1}x_{2}+x_{1}x_{3}+x_{2}x_{3})+P_{6}E=0.$
Finally, note that as $E$ can be written in terms of symmetric polynomial multiples of $x_{1},(x_{2}+x_{3})x_{1}$, this equation would be a relation between the first 5 generators of $Q_{0}(3,\mathbf{F}_{3})$. We have seen this is impossible, so we have $P_{6}=0$, and hence all of the $P_{i}$ must be 0.
Finally, note that the Hilbert series of the submodule of $Q_{0}(3,\mathbf{F}_{3})$ generated by these 6 polynomials is
$\frac{1+2t+2t^{2}+t^{3}}{(1-t)(1-t^{2})(1-t^{3})}=\frac{1}{(1-t)^{3}}.$
This is exactly the Hilbert series of $Q_{0}(3,\mathbf{F}_{3})$, so there are no more generators of $Q_{0}(3,\mathbf{F}_{3})$. ∎
Similar to how we only considered polynomials in the $(-1)$-eigenspace of $s_{12}$ for $\mathrm{sign-triv}$, we only consider generators in the $(-1)$-eigenspace of $s_{12}$ for $\mathrm{sign-triv-sign}$ and polynomials in the $1$-eigenspace of $s_{23}$ for $\mathrm{triv-sign}$. Note that this is sufficient to describe the roles of $\mathrm{sign-triv-sign},\mathrm{triv-sign}$, as both modules are generated by an element satisfying their respective constraints.
Lemma 6.6.
  1. 1)
    Let $T\in Q_{m}(3,\mathbf{F}_{3})$ generate a copy of $\mathrm{triv-sign}$. Then $T$ is the sum of a symmetric polynomial multiple of $E\prod_{i<j}(x_{i}-x_{j})^{2m}$ and a symmetric polynomial. Conversely, any symmetric polynomial multiple of $E\prod_{i<j}(x_{i}-x_{j})^{2m}$ generates a copy of $\mathrm{triv-sign}$ in $Q_{m}(3,\mathbf{F}_{3})$.
  2. 2)
    Let $T_{1}\in Q_{m}(3,\mathbf{F}_{3})$ generate a copy of $\mathrm{sign-triv-sign}$. Then $T_{1}$ is the sum of a symmetric polynomial multiple of $F\prod_{i<j}(x_{i}-x_{j})^{2m}$ and a symmetric polynomial multiple of $\prod_{i<j}(x_{i}-x_{j})^{2m+1}$. Conversely, any symmetric polynomial multiple of $F\prod_{i<j}(x_{i}-x_{j})^{2m}$ generates a copy of $\mathrm{sign-triv-sign}$ in $Q_{m}(3,\mathbf{F}_{3})$.
Proof.
1) We first prove the lemma for $m=0$. Consider some $T$ as above, and note that $(1-s_{12})T$ is contained in the sign representation, so by Proposition 2.9 we have $(1-s_{12})T=P(x_{1}-x_{2})(x_{1}-x_{3})(x_{2}-x_{3})$ for some symmetric polynomial $P$. Then note that $PE,T$ generate two copies of $\mathrm{triv-sign}$ with the same sign subrepresentation, so they generate a copy of
$(\mathrm{triv-sign}\oplus\mathrm{triv-sign})/\mathrm{sign}\cong\mathrm{triv-sign}\oplus\mathrm{triv}.$
So $T$ is the sum of $PE$ and a symmetric polynomial, as claimed.
Now, consider general $m$. By the above we have that any $T$ must be of the form $T=PE+Q$ for symmetric polynomials $P,Q$. Then since $T$ is $m$-quasi-invariant, we have $(1-s_{12})T=P(x_{1}-x_{2})(x_{1}-x_{3})(x_{2}-x_{3})$ is divisible by $(x_{1}-x_{2})^{2m+1}$. So $P$ is divisible by $(x_{1}-x_{2})^{2m}$, and it must also be divisible by $\prod_{i<j}(x_{i}-x_{j})^{2m}$ since it is symmetric.
The converse is clear.
2) This proof is similar to part 1). For $m=0$, any $T_{1}$ must have that $(1+s_{23})T_{1}$ is in a $\mathrm{triv-sign}$ representation, so $(1+s_{23})T_{1}=PE$ for some $P\in\mathbf{F}_{3}[x_{1},x_{2},x_{3}]^{S_{3}}$. Then $T_{1},PF$ generate a copy of
$(\mathrm{sign-triv-sign}\oplus\mathrm{sign-triv-sign})/\mathrm{triv-sign}\cong\mathrm{sign-triv-sign}\oplus\mathrm{sign},$
which implies the result for $m=0$. Then the extension to general $m$ is the same as in part 1). The converse is clear, as before. ∎
Finally, we can prove Theorem 1.3 for $p=3$. Note that this also implies Conjecture 5.2.
Theorem 6.7.
$Q_{m}(3,\mathbf{F}_{3})$ is freely generated by $1$, the two generators $K,L$ of $Q_{m}(3,\mathbf{F}_{3})_{\mathrm{sign-triv}}$ from Theorem 5.3, the two generators $K_{1},L_{1}$ from Proposition 6.1, and the generator $F\prod_{i<j}(x_{i}-x_{j})^{2m}$ from Lemma 6.6.
Proof.
Let us first show that there are no other generators of $Q_{m}(3,\mathbf{F}_{3})$. Assume for contradiction that there is some other generator $T$ of $Q_{m}(3,\mathbf{F}_{3})$. Then $T$ cannot generate a copy of $\mathrm{triv}$ by Proposition 2.9 and it cannot generate a copy of $\mathrm{sign-triv}$ or $\mathrm{triv-sign-triv}$ by Theorem 5.3 and Proposition 6.1. If it generates a copy of $\mathrm{sign}$, then by Proposition 2.9 it must be $\prod_{i<j}(x_{i}-x_{j})^{2m+1}$, but this polynomial is generated by $K,L$ by Lemma 4.4, so it cannot be a generator. If it generates a copy of $\mathrm{triv-sign}$, then it is $E\prod_{i<j}(x_{i}-x_{j})^{2m}$ by Lemma 6.6. But this is generated by $K_{1},L_{1}$ by Remark 6.3. Finally, by Lemma 6.6 the only generator that generates a copy of $\mathrm{sign-triv-sign}$ is $F\prod_{i<j}(x_{i}-x_{j})^{2m}$.
Finally, we show there are no relations between the 6 generators. Note that this also implies $F\prod_{i<j}(x_{i}-x_{j})^{2m}$ is a generator, since it is not generated by the other 5 generators. But this is clear: we already know there are no relations between the first 5 generators by Corollary 6.2. If there was a relation involving $F\prod_{i<j}(x_{i}-x_{j})^{2m}$, then note that since every generator is generated by the generators of $Q_{0}(3,\mathbf{F}_{3})$, this would induce a relation on those generators. Moreover, $F\prod_{i<j}(x_{i}-x_{j})^{2m}$ is the only generator not generated by the first 5 generators of $Q_{0}$, so the induced relation would be nontrivial. But there is no such relation by Proposition 6.5. ∎
Acknowledgements
We would like to thank the PRIMES program for making the project possible. This material is based upon work supported under the National Science Foundation Graduate Research Fellowship under Grant No. 2141064.
[…]